### v2

# Inverse design of potential metastructures inspired from Indian medieval architectural elements

Abstract
In this study, we immerse in the intricate world of patterns, examining the structural details of Indian medieval architecture for the discovery of motifs with great application potential from the mechanical metastructure perspective. The motifs that specifically engrossed us are derived from the tomb of I’timad-ud-Daula, situated in the city of Agra, close to the Taj Mahal. In an exploratory study, we designed nine interlaced metastructures inspired from the tomb’s motifs. We fabricated the metastructures using additive manufacturing and studied their vibration characteristics experimentally and numerically. We also investigated bandgap modulation with metallic inserts in honeycomb interlaced metastructures. The comprehensive study of these metastructure panels reveals their high performance in controlling elastic wave propagation and generating suitable frequency bandgaps, hence having potential applications as waveguides for noise and vibration control. Finally, we developed novel AI-based model trained on numerical datasets for inverse design of metastructures with a desired bandgap.
Keywords: Metamaterials, Metastructures, Interlaced metastructures, Hierarchical Metastructures, Fractals, Honeycomb, Elastic wave control, Artificial Intelligence, Multihead geometric attention, Multiscale residual network, Inverse design.
Introduction
Architectural intricate patterns have a long history, with certain examples dating back to the 4th century [1]. However, these early forms were the outcome of creative heuristic advancements of pattern generation in human civilization. It wasn’t until the late 19th century, when fundamental physics became better understood, that the development of systematic engineering principles made it possible to explore counter-intuitive physical characteristics of some patterns in periodic structures, generally known as metamaterials and metastructures. Today, it is often proposed as a panacea for controlling attenuation and wavefront divergence related to elastic and electromagnetic wave propagation and there exists a vast literature of which are based on meticulously designed lattice structures with special geometric patterns [2, 3, 4, 5, 6]. These engineered lattices may introduce exotic mechanical and electromagnetic properties, manipulating the behavior of electromagnetic waves [7], elastic waves as well as sound waves [8, 9, 10, 11]. By carefully designing these lattices, scientists can achieve extraordinary effects like negative stiffness [12], negative mass [13], and negative Poisson’s ratio [14]. The study of lattice geometries in metamaterials opens a pathway to myriad innovative applications ranging from optics to acoustics, wearable devices [15, 16], advanced space structures [17, 18], next-gen automobiles [19, 20], robotic manipulators [21], bionic grippers, [22] and anti-seismic structures [23, 24]. These metamaterials exhibit characteristics such as high specific stiffness [25, 26, 27, 28], negative Poisson’s ratio [29, 30], negative thermal expansion coefficient [31], negative reflection [32, 33], wave alteration [34], sub-wavelength bandgap [35, 36] and distinctive electro-dynamic phenomena [37, 38]. The generation of a large range of novel periodic structures at different length scales has ushered in phenomenal developments of metamaterials.
Figure 1: The Tomb of Itmad-ud-Daula is a Mughal mausoleum in the city of Agra in the Indian state of Uttar Pradesh (27.1929 N, 78.0310 E). The architecture of this monument is unique in terms of crafting which uses a combination of pietra-dura and calligraphy-based decorative art techniques. It is considered to be the pioneer of the creation of funerary tombs in Mughal-India which was subsequently magnified in the Taj Mahal. However, with its marble latticework and exquisite carvings, the tomb is considered as a work of art even more delicate than the Taj Mahal. (b to d). Consequently, the latticework is depicted through the carved panels of calligraphic design. The walls, floors, and niches are decorated with beautiful geometrical patterns. (e and f). Here, we have specifically focused on the Jali screens at the centre of the niches all around the tomb. They were used extensively during the Mughal reign for elegance and their ability to facilitate the entry of natural light and air inflow. Often it is also used as part of a passive cooling system. (g) shows the corresponding structural models (LS1-LS9).
Figure 2: Different types of metastructure unit cells can be categorized based on their lattice structures. (a) Five popular unit cells are used as primary lattices. (b) A complex array of primary lattices can be used to create different possible interesting metastructure unit cells. (c) Multi-unit patterns represent the arrangement of unit cells to form complete 2D and 3D metastructures.
Mechanical properties of simple metastructures such as honeycombs with positive and negative re-entrant angles have been extensively studied through experimental avenues, numerical simulations, and mathematical analyses [39, 40, 41, 42, 43, 44]. In recent years, there has been a growing interest in developing engineered materials using a range of 3D and 4D additive manufacturing techniques [45, 46]. Recent advancements in precise additive manufacturing techniques at the microscale have further enabled the fabrication of new lattices with complex topology and interlaced geometry, which are challenging to achieve using traditional subtractive manufacturing methods. Such metastructures have distinct behaviors when subjected to external stimulus such as mechanical excitation [47, 48, 49, 50, 51, 52, 53].
Geometric design and pattern have played a significant role in the spiritual traditions of most of the religions. For instance, concentric circles and triangles are widely used as sacred patterns in the Hindu and Buddhist ‘Mandala’; while rectangles and squares are preferred in the ‘Feng Shui’ design of Chinese spiritual tradition. The construction of medieval European cathedrals was often based on the ‘Latin Cross’. Fractal geometries were the most popular design used in the construction of various Hindu temples. The pinnacle of embedding geometric motifs is found in ‘Islamic Geometric Patterns’ which involve more complicated shapes such as hexagons, octagons, decagons, and multi-point stars [54]. One of the interesting applications of such motifs in tessellated form is called ‘Mashrabiya’ which is known to be originated in Baghdad during the 12th-century Abbasid period. Its interlaced arrangements, which emerged from simplicity and unity of natural as well as geometric motifs, exhibit some fundamental relationships that become in-turn hierarchical or fractal in complex architectures. The architectural applications of ‘Mashrabiya’ had rapidly developed in the medieval Persian and Indian civilizations due to the influence of Islamic religion and mysticism. Such patterns, popularly known as ‘Jali’ in India, were crafted not only for aesthetic purposes but also for their utility in terms of transmitting indirect sunlight, capturing fresh air, passive cooling, maintaining communication with privacy, etc. (some of the patterns are shown in Fig. 1(b-f)). The first instance of developing such patterns in India was found in the Alai Darwaza in the Qutb Minar complex of Old Delhi. The tomb of Itimaduddaula, a magnificent testament to the Mughal style enriches Agra’s vibrant tapestry. Considered as the inspiration for the Taj Mahal, the tomb also known as the ‘Baby Taj Mahal’, is a stunning example of Mughal architecture (Fig. 1 (a)). This mausoleum displays a wonderful variety of geometric arabesques. The strong influence of the Persian style coupled with numerous Indian motifs is visible in this structure. The related lattice structures can be classified based on different unit cell types, considering angle, arrangement, and symmetry (Figure 2(a)). We investigate various interlacing arrangements of these unit cells (Figure 2(b)) and demonstrate how different arrangements can be used to generate complete metastructures (Figure 2(c)). These metastructures provide a foundation for further studies evaluating their performance in controlling elastic wave propagation and vibration. Further, we have developed an AI-based forward analysis model to predict the transmissibility of a metastructure. Also, an AI-based inverse design of metastructure is developed to design new lattices for the desired bandgap creating end-less possibilities of generating suitable patterns focused to an application [55, 56, 57, 58, 59, 60].
Figure 3: Schematic diagram illustrates latticed mechanical metastructures featuring different interlaced lattice structures (LS1-LS9). Dynamic analysis of the metastructures is performed and mode shapes of lattices at a frequency corresponding to minimum force transmissibility for in-plane excitation are shown. The attenuations corresponding to each geometry are obtained by using the transmissibility plot for each structure. We have highlighted distinct attenuation zones within the transmissibility response of metastructures through green-shaded patches.
Figure 4: Conducting a comparative analysis, this study investigates the impact of scaling (x, 1.5x, 2x) of the unit cell on the bandgap, with the resulting BG ratio illustrated in a bar diagram. The basic building blocks of unit and sub-unit cells are represented in the middle column of the figure. Subunit cells are used as an interlaced component for the primary structure from which each configuration is constructed. The patterns of latticed mechanical metastructures are 3D Printed using PLA having different primary lattice structures namely (a) Honeycomb-honeycomb interlaced (LS1) (b) Honeycomb interlaced with honeycomb pattern (LS2) (c) Honeycomb interlaced with a star (LS3) (d) Chiral interlaced structure (LS4) (e) Ring interlaced structure (LS5) (f) Fishtail structure (LS6) (g) Star-honeycomb structure (LS7) (h) Honeycomb-auxetic structure (LS8) (i) Honeycomb-auxetic interlaced structure (LS9).
Figure 5: Experimental validation of computational result with variations in mass inserts pattern is shown, and the corresponding bandgaps are marked by green patches. (a) Simple interlaced honeycomb structure (LS1) (b) Diamond pattern with a broken symmetry at the middle (c) Cross pattern (d) Forward arrow pattern (e) Diamond pattern (f) Top and bottom row are filled middle row is empty (g) Top and bottom rows are filled with mass inserts while the middle row has two-gap arrangement with no inserts (h) Upward pointing arrow pattern (i) Top and bottom row inserts are arranged alternatively while middle row has two-gap arrangement with no inserts.
Results
Dynamic characterization and determination of the Bandgap
Selected latticed metastructures (represented in Fig. 3), generated based on architectural elements (described in Fig. 2) have an overall size of 204 mm x 58 mm and a thickness of 5 mm. The size was carefully chosen keeping in mind the manufacturing constraints as well as the requirement of minimum number of periodic units (greater than 7 units) in the tessellated structure to express the metamaterial behavior [13]. On both sides, 15 mm solid frames are appended with the latticed structure in the longitudinal direction (see Supplementary Fig. S1) for ease of clamping and excitation during dynamic testing. For each latticed structure, there is a frequency band in which the structure may work as an attenuator signifying the absence of wave propagation, and the same is highlighted in Fig. 3. In the middle column (Fig. 3), we have shown the related mode shapes obtained using Ansys corresponding to minimum force transmissibility within the attenuation zone. Different structures have different damping ratios depending on the lattice geometry. The damping ratio is calculated by the half-power bandwidth method (highest for LS1 and the lowest for LS9), which is shown in the supplementary Fig. S11.
Latticed metastructures with different sub-lattice structures are shown in Fig. 4. Figure 4(a-c) illustrates the primary unit cells as honeycomb sub-lattice, single honeycomb, circular pattern of honeycomb cells and star structure respectively. In Figure 4(d), the primary unit cell has a chiral structure with sub-lattices exhibiting a spider web-like pattern. Furthermore, Fig. 4(e) features a primary unit cell as a ring structure interlaced with other ring structures. Figure 4(f) is a fishtail structure. Fig. 4(g) shows the primary unit cell as a series of honeycombs arranged to form a star at the center of the structure. Figure 4(h, i) are combinations of honeycomb and auxetic structure. In the former, these structures are present in a series combination, while in the latter, the auxetic structure is interlaced with the honeycomb structure. The dimensions of the unit cells are detailed in Supplementary Table S1. After creating all the metastructures of specified dimensions, we scaled them up by 1.5 and 2 respectively to capture the scale effect of unit cells in the wave transmission pattern. The scaling effect is illustrated with the help of the Band Gap (BG) ratio in the bar diagram (Fig. 4).
Performance measurement of lattice with resonators
In figure 5, we have selected a primary honeycomb lattice structure interlaced with a single honeycomb unit (LS1) at the centre. Our objective is to study the effect of mass insertion in the central honeycombs on the bandgap characteristics of the metastructure. The metastructure is chosen as per the ease of mass insertion and enables us to observe changes in the attenuation band, which are influenced by local resonances occurring at frequencies that vary with the insertion location. Metallic hexagonal inserts are made up of structural steel (as shown in Supplementary Fig. S5), and their detailed properties and composition are shown in Supplementary table S3. Given that the lattices are fabricated through 3D printing using Polylactic Acid (PLA), and subsequently metallic inserts of much heavier density are used inside the lattice, these inserts function as local resonators due to the density difference between PLA and steel. These characteristics make the metastructure well-suited for vibration control applications. The presence of inserts significantly influences the wave transmission patterns. Furthermore, the strategic placement of the inserts are considered in such a manner that the inserts themselves make a pattern inside the lattice structure. By changing their location, we can control the bandgap for planar excitation in 1D. In Figure 5(a), we present the effect of random vibration through the simple lattice, while in Fig. 5(b) we explore the effect of wave transmission in broken symmetry of diamond pattern where the middle row of the lattice is fully filled while the top and bottom row is filled with metallic inserts with a definite pattern. In Fig. 5(c), the effect of wave transmission in the cross-shape pattern is shown where the inserts are arranged alternatively in the middle row of lattice and the top and bottom row is completely filled. In Fig. 5(d) lattices are filled in such a manner that it forms an arrow indicating towards the right direction. Figure 5(e), represents the diamond pattern where the middle row is fully filled; however, the top and bottom rows are alternatively filled. In Fig. 5(f), the top and bottom rows are fully filled and the middle row is fully empty. In Fig. 5(g), the top and bottom rows are fully filled; however, the middle row is filled such that there is a spacing of two units between the consecutive inserts. In Fig. 5(h), the top row is alternatively filled starting with a blank cell while the bottom row is also alternatively filled starting with a filled cell, and the middle row is fully filled except for the first and the last cell representing upward pointing arrow. In Fig. 5(i), the top and bottom rows are alternatively filled while the middle row is filled with two-unit spacing. Through the study of dynamic response corresponding to harmonic excitation for all these combinations, we aim to find out which pattern location is more sensitive and affects more at a specific frequency range of wave transmission. We observed that generally, the middle row exhibits higher sensitivity compared to the top and bottom rows. Additionally, by arranging different combinations of inserts, we achieved up to 48 % bandgap enlargement, multiple bandgaps, and the ability to shift the bandgap across the desired frequency range. Further, the experiments corresponding to each iteration are carried out to validate our simulation results by using a Laser Doppler Vibrometer (LDV). Random vibration up to 10 KHz is performed using LDS shaker with the gain set as unity. Detailed experimental setup can be seen in Supplementary Fig. S10. As may be observed in Fig. 5, our experimental results show good agreement with the simulation. The printing pattern, and infill density (presented in Supplementary Table S2), also affect the attenuation band.
Bandgap characteristics of honeycomb based interlaced structures
Figure 6: (a) A comparison of bandgap at different frequency levels of different structures interlaced in the honeycomb structure. (b), (d) The comparison of static performance is obtained through analytical computations of different structures. The displacement control mechanism is set as 5 mm compression of initial height of samples (c) Comparison of bandgap in different order (n=0,1,2) of the fractal structure.
Comparative analysis of the effect of different interlaced structures on bandgap shift (Fig. 6(a)) is carried out in various frequency ranges. The band gap occurs at a frequency band comparatively higher than that of the simple honeycomb structure. Notably, for the LS3 structure bifurcation in bandgap occurs, and for the last structure (LS9) multiple bandgaps occur both for the higher and lower frequency ranges. Furthermore, we have compared the bandgap range of various interlaced honeycomb unit cells (Fig. 6b) along with the force-deflection relationships. The force-deflection relationships are linear for all the patterns except for LS9.
Figure 6(c) illustrates the fractal structure of different order (n=0,1,2) and their corresponding bandgap shifts. In different structures, as the fractal order increases, the bandgap shifts to higher frequency ranges, and for further increase in fractal order multiple bandgaps occur. Furthermore, on the right side we have represented the force-deflection relationship for the fractal structure (see Fig. 6d). With the increase in the fractal order, the structure becomes stiffer, with 196.32% increase from zeroth to first-order fractal and 33.61% increase from first to second-order fractal. From this study, we can conclude that with increasing the structural complexity inside the structure, the bandgap starts to shift towards the higher frequency range, and for higher-order fractals, it also bifurcates which is verified by taking different fractal structures inside the honeycomb and explained in the Supplementary Fig. S8.
AI-driven Inverse Design of Metastructures
With the phenomenal expansion potential of the geometry of metastructures, it is possible to use AI efficiently for both forward analysis in predicting the bandgaps as well as inverse design where for a given bandgap a suitable pattern may be obtained. The complete flow diagram of a possible AI-based inverse design of metastructures is depicted in Fig. 7. The entire method comprises two components: (i) the forward analysis model and (ii) the inverse design model. The forward analysis model acts as a surrogate model for training the inverse design model. It is trained independently for the forward analysis of the metastructures with known transmissibility.
Figure 7: AI-based inverse design of metastructures: (a) Training strategy of forward analysis model to predict the transmissibility for a given metastructure (b) Training strategy of inverse design model to predict the metastructure for given transmissibility spectrum.
Forward Analysis Model
The forward analysis model consists of repeated blocks of geometric feature extraction modules capable of learning complex metastructure geometry having primary, secondary, and tertiary complexity. Each geometrical feature extraction module consists of a multi-scale geometric attention mechanism having multiple heads of FEM-inspired spatial attention (FSA) blocks. The FSA is designed to learn the dependencies of adjacent connections at different scales using different kernel sizes of convolutional modules. Each FSA consists of an input convolutional layer for feature transformation followed by a spatial attention block [61, 62]. The complete block diagram of the forward analysis block, multi-scale GA module, and the FSA block are provided in the supplementary Figures S13 and S14. The forward analysis model was implemented with 12 layers of geometric feature extraction blocks and 8 heads of FSA in each block. The kernel sizes of the input convolutional layer in 8 FSA heads are selected as $(1,1,1,3,3,3,5,5)$. This combination of kernel sizes puts more attention on the smallest structures (tertiary), than the secondary, and the primary structures. The output of the geometric features extraction module is processed by a regression head that transforms the geometric features to the transmissibility of dimension $(1000,1)$. The detailed block diagram and the descriptions are provided in the supplementary file: Fig. S15 and Section E.1.
Inverse Design Model
The inverse design model is developed to obtain the metastructure geometry using transmissibility plots for the desired bandgap. The detailed block diagram of the inverse design model is provided with the supplementary materials (Figures S16 and S17). The key contributions of the inverse design model are the conversion of transmissibility into latent space using a multi-head 1-D self-attention mechanism, followed by constructing the interlaced patterns using a multiscale residual network. The 1-D self-attention mechanism captures the sequential dependencies of the transmissibility spectrum, [63] and generates a rectified latent space of shape $(3\times 128\times 256)$. The latent space is further extended to a new higher space using reflection padding and then non-linearly down-sampled back to the same dimension using multiple consecutive convolutional layers with ReLU activation.
Now the transformed latent space is passed through a multiscale residual network having three parallel convolutional paths with three different kernel sizes: $(1\times 1)$, $(3\times 3)$, and $(5\times 5)$ as shown in Fig. S17(b). These sizes have been selected based on the institution to transform the latent space into the three levels of interlaced structural geometry. The smallest kernel is responsible for learning the tertiary structures, the medium size helps to learn the secondary structural geometry. Similarly, the largest value $(5\times 5)$ is responsible for learning the primary geometry. The final output of the multi-scale residual network is the interlaced patterns. It is now upsampled using three layers of 2-D transposed convolution and then transformed non-linearly using 2-D convolution with hyperbolic tangent (Tanh) activation function to produce the final geometry.
Training
For the training of the proposed AI-based inverse design of the metastructures, a dataset having 720 samples of different metastructures is prepared using different interlaced structures as shown in supplementary Fig. S12. In the unit cell of these structures, the primary structure is a honeycomb and the secondary structures are different combinations of honeycomb, star, circle, hierarchical honeycomb, and auxetic interlaced structure. Then the corresponding transmissibility is calculated using FEM analysis.
The samples are divided into two portions: training and validation with a ratio of 90:10. The AI-based inverse design mechanism is trained in two phases.
  1. i)
    Forward Analysis Model: In the first phase, the forward analysis model is trained using the training samples to predict the transmissibility of a given metastructures. The training loop is depicted in Fig. 7(a). The comparison of the predicted transmissibility obtained by the forward analysis model and the true transmissibility obtained by the FEM analysis are compared in Fig. 8.
  2. ii)
    Inverse Design Model: Once the forward analysis model is fully trained, it is used as a surrogate model for the training of the inverse design model as displayed in Fig. 7(b). The loss function $(\mathcal{L})$ for the weight update of the inverse design model is constructed as provided in equation (1)
    $\mathcal{L}=\mathcal{L}_{1}\;+\;\beta_{1}*\mathcal{L}_{2}\;+\;\beta_{2}*\mathcal{L}_{3}$ (1)
    where, $\beta_{1}$ and $\beta_{2}$ are the penalty terms for the inclusion of losses $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$. The values of $\beta_{1}$ and $\beta_{2}$ were selected as 0.1 and 0.6, respectively.
    $\mathcal{L}_{1}\;=\;\text{MSE}$: mean square error between predicted transmissibility on generated metastructure and the transmissibility obtained from the FEM method on original metastructure.
    $\mathcal{L}_{2}\;=\;1-\text{SSIM}$: structural similarity index measure between the generated metastructure and the original metastructure [64].
    $\mathcal{L}_{3}\;=\;\text{MMD}$: maximum mean discrepancy between the generated metastructure and the original metastructure (MMD) [65].
    Further details of these losses are provided in the supplementary material, Section E.3.
Validation of AI-model for metastructure design
The metastructure produced by the inverse design model contains colored noises due to training with a small value of batch size (small batch size was chosen due to the limited number of samples in the dataset). Therefore, a noise removal filter is applied to produce a clear image of the obtained metastructure during validation. The algorithm for the noise removal is provided in supplementary material, Section F. The AI-based inverse design of metastructures is validated under two cases:
Figure 8: Validation results for AI-based inverse design of metastructure. (a) Original metastructure of the test sample. (b) Input transmissibility obtained by FEM for the original metastructure. (c) Metastructure predicted (generated) by the inverse design model using the input transmissibility. SSIM = 0.991, closer to unity validates the exact match with the original image. (d) Comparing the transmissibilities of the original metastructure obtained from the forward model and FEM. (e) Comparing the transmissibilities of the generated metastructure obtained from the forward model and FEM.
Validation on known metastructure as test sample:
The comparison of generated metastructure for one test sample from the test data (10% of available 720 samples) is provided in Fig. 8. Figures 8(a), (b) show the test metastructure and its transmissibility obtained by the FEM method which is assumed to be true (input) transmissibility for inverse design. The predicted metastructure by the inverse design model exactly matches the original metastructure with an SSIM score of 0.991 as shown in Figure 8(c). Forward analysis is performed on both the original and the predicted metastructures. The predicted transmissibilities are compared with the input (true) transmissibility in Fig. 8(d) and (e). It can be observed that MSEs for transmissibility obtained from the original metastructure and the predicted metastructure are 0.0834 and 0.1939 respectively. The MSE for predicted transmissibility by forward analysis on the predicted metastructure is comparatively larger because of the existence of noises in the predicted metastructure.
Validation on Artificial Transmissibility:
The AI-based inverse design model is further validated for practical deployment to design a metastructure facilitating the formation of a desired frequency bandgap. Hence, an artificial transmissibility is created having the desired bandgap and fed to the inverse design model. The generated metastructure and the corresponding transmissibilities obtained by the forward analysis model and FEM are shown in Fig. 9. The obtained metastructure is the combination of a honeycomb and hierarchical honeycomb substructures. The MSE between the predicted transmissibility and the input transmissibility is 0.7408.
Figure 9: AI-based inverse design of metastructure for a desired bandgap. (a) Artificial (input) transmissibility having the desired bandgap as highlighted. (b) The generated metastructure obtained from the inverse design model, which facilitates the desired bandgap. (c) Validation of the generated metastructure design by comparing the transmissibilities obtained by the forward analysis model and FEM.
Discussion
It has been found experimentally and analytically from the previous study that a simple lattice structure has the ability to block waves for a certain frequency range when it passes longitudinally throughout the lattice. Inspired from the architectural patterns, we have designed mechanical metamaterial lattices with diverse interlaced structures. A comprehensive study has been performed through simulation-based analysis using Ansys. Experiments have been conducted to validate the transmissibility property corresponding to different lattice configurations. This work shows the blueprint of metamaterials, including the fact that the effective properties of the lattice are contingent on collective patterns and subpatterns of the unit cells. Furthermore, the tunability of the bandgaps through design resulted in larger and multiple bandgaps occurring in Frequency Response Function (FRF) plots. The reported band gap caused by the lattice structures can be further modified by adding metallic inserts with different combinations of insert patterns. These metallic inserts interact with the geometry and due to the large density difference as compared to lattice material (PLA), they act as local resonators, providing flexibility to adjust bandgaps. Some significant negative peaks are obtained from the transmissibility plots that signify wave attenuation and show good agreement with the simulation results, as shown in Fig. 5. The discrepancy in some amplitude and peak are due to nonconsideration of damping in simulation. Apart from damping, printing speed, orientation, infill pattern, and infill density of 3D printing are also affected as discussed and shown in Supplementary Table S2. These findings unveil a deep connection between lattices and band gaps and have broad potential applications in the field of vibration isolation for high-speed trains and in serving as surface waveguides and absorbers.
The second part of this study on the inverse design of metastructure using geometric attention-based deep architecture is a benchmark in the field of inverse design of complex structure whether it is metastructure for the desired mechanical response or any other complex pattern for desired physical properties. The major challenge in the training of an AI-based inverse design model is the availability of sufficient number of training samples. The higher the complexity of the structure, the higher the number of samples required for the training. Our proposed model uses FEM to learn the dependencies of adjacent connections in the geometrical structure. Therefore, it can easily extract the information embedded in the complex structure and learn the relationship between the input geometry and the output transmissibility. Further, the multiscale geometrical attention mechanism provides the capability to learn the multiple levels of complexity. With the same concept, multiscale feature transformation using the multiscale residual network enables the inverse design model to reconstruct multiple levels of geometry, like a honeycomb and hierarchical honeycomb substructures. Therefore, the proposed framework of the geometrical attention-based feature learning and the inverse design can be useful for any other scientific design problem.
In summary, our paper reflects the following novelties:
  • Discovery of novel functionality in a set of ancient patterns from the mechanical metastructure perspective such as vibration transmission and attenuation.
  • Identification of two unique parameters: interlaced metastructure and fractal order in such patterns followed by their application in 3D printed 2D mechanical metastructure panels.
  • Further modifications of such metastructures by including local resonators in the form of mass inserts to improve bandgap characteristics through experimental studies.
  • Creation of novel AI-based forward analysis model using multiscale geometric attention mechanism which can predict the transmissibility as well as bandgap of a metastructure subjected to base excitation.
[…]
  • Further, we have been able to create complex multiscale patterns successfully and predict their bandgap characteristics through the transmissibility plot using the current AI model. We believe this will facilitate the generation of many customizable complex patterns for mechanical metastructures efficiently.
  • A novel inverse design model of metastructure based on NLP inspired multiheaded attention mechanism with 1D position encoding and multiscale residual network has also been demonstrated by providing artificial transmissibility for a desired frequency bandgap. The inverse design model is found to be much more efficient to reconstruct the complex metastructure with similarity index of 0.991.
Conclusions
This research study highlights the translational application potential of vast range of patterns and motifs developed in the field of architecture into the metamaterial-based wave attenuation. Thus, the utilitarian values of aesthetic constructs in the field of vibration and sound engineering may open up new systems for wave control including bandwidth control, subwavelength acoustic imaging, and generation of acoustic black holes. Additionally, a systematic investigation process has been created on the basis of root lattice identification, multi-scale lattice formation through a hierarchic system, development of hybrid lattice structure, and the effect of interlacing in the vibration mitigation potential initiated from the historic geometric patterns found in the tomb of I’timad-ud-Daulla. In general, the simple lattice structures typically act as attenuators, however, this study demonstrates that the interlaced structure offers an enhanced vibration attenuation property depending on the interlace geometry with improved control over the attenuation frequency range. Such effects can be further enhanced, in the presence of and through strategic positioning of the metallic inserts inside the lattice structure which may result in significant expansion of the bandwidth. This method has been enhanced with the help of AI-based inverse design models to predict metastructures that exhibit the desired bandgap and the corresponding transmissibility. Although this study concentrates on predominantly honeycomb-shaped interlaced structures, the principle of incorporating mass resonators can be applied to other geometries to manipulate bandgap frequencies. Moreover, the concept may be extended to three-dimensional interlaced structures with diverse lattice and interlace configurations across multiple length scales. Another possibility of future expansion is related to the insertion of smart materials in the lattice structure such that the stiffness and the damping can be controlled over time; thus, transforming it into 4-D metamaterial.
Methods
Experimental testing
The study utilizes a non-contact technique to measure vibration to assess the response of different latticed metastructures. Nine 3D printed samples, each showing unique metastructure, are subjected to dynamic testing using a 3D Laser Doppler vibrometer (LDV) from Polytec (Details of the experimental setup presented in Supplementary Fig. S10). These samples are attached to an Electrodynamic shaker system in cantilever mode. The laser head of LDV captures surface velocity by analyzing the phase shift between the incident and reflected laser beams. Data acquisition and signal processing are executed through the NI-DAQ system. The base excitation technique is utilized to determine transmissibility. In this method, the LDS shaker generates 3200 Fast Fourier Transform (FFT) lines, and the pseudo-random excitation with a frequency range of 0 to 10 KHz. Displacements of the top and bottom surfaces of the samples are recorded to calculate transmissibility after post-processing of the data.
Finite element simulations
Finite element analysis-based simulation has been carried out using Ansys, employing the Ansys Parametric Design Language (APDL) solver for all the nine structures, and the host material is chosen as PLA (detailed properties given in Supplementary Table S3). The entire structure was discretized by using tetrahedron elements with a 2 mm element size, to ensure convergence for interlaced geometries. The smaller edges of the base structure were fixed, while the longer edges remained free. A dynamic force of amplitude 10 N was applied to the base, and the response was calculated on the opposite phase. Modal analysis was performed to identify all modes up to 10 KHz frequency, followed by harmonic analysis to obtain the transmissibility. The mode superposition method was used to link the modal analysis results with the harmonic analysis input. Frequency response functions (FRFs) were generated over the frequency range of 0 to 10 KHz with a 10 Hz frequency resolution; this step size is preferable to accurately capture the response. A structural damping ratio of 0.001 was incorporated to account for the viscoelastic effect of PLA substrate.
Data and code availability
All relevant data and codes will be available upon reasonable request from the contact author. The research data are presented either in the main file or in a supplementary file.
Acknowledgements
The authors want to thank Mr. Gyanendra Prakesh Tripathi, who provided insight to conduct the experiments, and Mr. Abhishek Kumar Singh, who helped to 3D print the samples. The photo credits for Figure 1(a-f) are given to Chitralekha Bhattacharya and Bipasana Bhattacharya.
Author information
Authors and Affiliations
Department of Mechanical Engineering, Indian Institute of Technology Kanpur, India
Bishakh Bhattacharya, Tanuj Gupta, Arun Kumar Sharma, Ankur Dwivedi, Subhadeep Sahana, Ashish Awasthi
Department of Precision and Microsystems Engineering, Delft University of Technology, Netherland
Vivek Gupta
Department of Mechanical Engineering, Motilal Nehru National Institute of Technology Allahabad, India
Suryansh Pathak
Author contribution
Bishakh Bhattacharya: Conceived the primary idea, reviewed the manuscript, critically examined the results, checked the data, wrote the manuscript, and acquired funding. Tanuj Gupta: Expanded the idea jointly, modeled all the structures, and designed and performed numerical simulations. Fabricated the samples and carried out related experiments and material testing, documented the results, reviewed the manuscript, critically examined the results, and checked the data, wrote the manuscript. Arun Kumar Sharma: Concived the idea of AI-based inverse design of metastructures, and implemented a new strategy of feature learning using multiscale Geometric Attention and the idea of reproducing the geometry with multiple levels of complexity using the multiscale residual network. Code Implementation, training, and validation of the AI models and writing the AI-based inverse design in the manuscript. Critically examined the manuscript. Ankur Dwivedi: Conceived and expanded the idea jointly in the inverse design of metastructures, critically examined the results, checked the data, wrote and reviewed the manuscript. Vivek Gupta: Expanded the idea jointly, reviewed the manuscript, critically examined the results, checked the data, and wrote the manuscript. Subhadeep Sahana: Carried out numerical simulations to generate synthetic data for training the AI model and critically analyzed the paper for logical conclusions. Suryansh Pathak: Conducted the experiments, performed numerical simulations, and checked the data. Ashish Awasthi: Carried out numerical simulations to generate synthetic data for training the AI model.
Competing interest
The authors declare no competing interests.
Supplementary Material
Appendix 1 Supplementary Methods
A Sample Fabrication
Figure S1: This figure represents a typical interlaced self-similar lattice metastructure employed for sample fabrication. The important geometric parameters are prominently displayed, alongside the designated active lattice region. This intricate design serves as the blueprint for all samples of our study. Some of the sublayer details are shown in Figures S2 and S3.
A.1 Structural design
The modeling of all lattice structures is carried out using SolidWorks modeling software. For all the structures, the length ($l$) is taken as 174 mm, width ($w$) 58 mm, and thickness ($t$) 5 mm respectively. On both sides of the structure, an extension of a 15 mm solid structure is also included for holding and testing purposes as shown in Fig. S1. The final dimension of the latticed structure ($L\times b\times t$) is $204\times 58\times 5$ mm which is uniform for all the structures so that the results will not get affected by the size of the samples [66]. Each sample unit cell has an interlaced structure. Some of them are interlaced inside a honeycomb structure, while others are interlaced with different structures like ring or fishtail patterns. The unit cell of each structure is shown in Fig. S2 and Fig. S3 and their detailed dimensions are provided in Table S1.
Figure S2: Intricate detail of unit cell for each lattice structure (LS1 - LS5). Precise dimensions are provided in table S1 which govern the macroscopic properties of the materials.
Figure S3: Intricate detail of unit cell for each lattice structure (LS6 - LS9). Precise dimensions are provided in table S1 which govern the macroscopic properties of the materials.
Table S1: Detailed dimensions of structures (LS1-LS9) represented in Figures S2, S3.
Parameter Value Parameter Value Parameter Value
LS1 $l_{1}$ 21 mm LS3 $d_{2}$ 2.33 mm LS6 $r_{2}$ 4 mm
$t_{1}$ 2 mm LS4 $l_{1}$ 12.89 mm LS7 $l$ 5.1962 mm
$l_{2}$ 8 mm $l_{2}$ 3.71 mm $t$ 1.5 mm
$t_{2}$ 1 mm $l_{3}$ 2.71 mm $b$ 0.8 mm
LS2 $l$ 21 mm $t_{1}$ 2 mm LS8 $l_{1}$ 18 mm
$t_{1}$ 2 mm $t_{2}$ 1 mm $l_{2}$ 16.17 mm
$t_{2}$ 1 mm $b$ 6.075 mm $t$ 2 mm
LS3 $l_{1}$ 21 mm $r$ 5.077 mm LS9 $l$ 21 mm
$t_{1}$ 2 mm LS5 $r$ 8.5 mm $t$ 2 mm
$\alpha$ $16.28^{\circ}$ $t$ 2 mm $\alpha$ $80^{\circ}$
$d_{1}$ 5.5 mm LS6 $r_{1}$ 6 mm
A.2 Additive manufacturing
Figure S4: (a) Additive manufacturing: Creating intricate three-dimensional samples layer by layer. (b) Finished 3D printed samples.
Metamaterial structures were modeled using SolidWorks 2022 and subsequently fabricated using a 3D printer that worked on a fused deposition method and has a printing accuracy of $\pm$ 0.1 mm. This technology prints samples layer by layer. The opted 3D printing specifications included a layer thickness of 0.2 mm and a $50\%$ infill density of the pattern. Due to the limitations of 3D printing technology, it was observed that the mechanical properties of the printed material were affected by factors such as layer orientation, infill density, and infill pattern [67, 68]. To maintain consistency, all sample structures were printed using the same specifications as shown in Table S2, and the corresponding samples are shown in Fig. S4. The 3D printing material PLA from eSun was utilized to print the sample. The filament diameter of the PLA is 1.75 mm. The mechanical properties of PLA sample, detailed in Table S3 include density = 1240 $Kg/m^{3}$, Poisson’s ratio = 0.33, and Modulus of Elasticity = 3.5 $GPa$.
Table S2: A detailed specification of 3D printing process while printing samples.
Sr. No. Specification in 3D printing Values
1 3D printing material PLA of 1.75 mm diameter
2 Layer thickness 0.2 mm
3 Infill pattern Zig-zag
4 Infill density $50\%$
5 Printing speed $50$ mm/sec
6 Print temperature $210^{\circ}C$
7 Temperature of build plate $60^{\circ}C$
8 Top and bottom layer thickness 0.8 mm
A.3 Metallic insert fabrication
Figure S5: (a) Mild steel hexagon inserts cut from plate by wire EDM process which is used as a resonator within the lattice structure. (b) Assembly of inserts within the lattice structure.
Metallic inserts, when strategically integrated into the matrix of a mechanical metamaterial, may function as local resonators. These components play a pivotal role in manipulating the attenuation bandwidth, enabling the design of structures with extraordinary characteristics unattainable through conventional materials.
Mild steel hexagonal shape inserts have been chosen for the selected interlaced honeycomb lattice. Initially, hexagon modeling is done on SolidWorks 2022 and exported as a DFX file for the wire EDM cutting using Ezeecut NXG wire EDM, which has a tolerance of $\pm$ 0.002 mm and a best surface finish of 1.5-2 $\mu m$. These hexagon shapes are then cut from a 5 mm thick plate, as shown in Fig. S5, resulting in inserts with an average weight of 1.0693 grams with a standard deviation of 0.0022 grams. The inserts function as a local resonator within the lattice structure due to their large density difference with PLA. The mechanical properties of the mild steel sample, detailed in Table S3, include density of 7850 $Kg/m^{3}$, Poisson’s ratio of 0.33, and Modulus of Elasticity of 200 GPa.
B Material Testing
B.1 X-Ray fluorescence spectrometry (XRF)
Figure S6: X-ray fluorescence (XRF) spectroscopy setup for identifying the elements in mild steel sample.
Table S3: Mechanical properties of PLA and Mild Steel, where subscript $p$ for PLA and $m$ for Mild Steel.
Material Property Value Material Property Value
$E_{p}$ 3.5 $GPa$ $E_{m}$ 200 $GPa$
$\nu_{p}$ 0.33 $\nu_{m}$ 0.33
$\rho_{p}$ 1240 $Kg/m^{3}$ $\rho_{m}$ 7850 $Kg/m^{3}$
The main objective of XRF is to identify and quantify the elements present in the material (stainless steel hexagonal inserts) without destroying them. A $25\times 25\times 5$ mm stainless steel sample of metal plate is prepared and sample is typically cleaned to remove surface contaminants and sample is irradiated with high-energy X-rays. The atoms in the sample emit characteristic X-rays in response to the excitation, and the emitted X-rays are detected and analyzed to determine the elements present and their concentrations (as shown in Fig. S6). The details of the element present in the sample are presented in Table S4.
Table S4: Component analysis of mild steel by X-ray fluorescence spectroscopy (XRF)
[…]
1 Fe 99.4
2 Mn 0.321
3 Si 0.128
4 Al 0.115
5 Cl 0.0491
6 S 0.0220
7 Ca 0.0141
B.2 Mechanical characterization: Dynamic mechanical behavior of PLA
Figure S7: Result of mechanical analysis. (a) Variation of damping parameter with temperature for PLA at various frequency ranges. (b) Storage modulus of 3D printed structure.
The DMA7100 is used to analyze the PLA sample (with zigzag infill pattern) of standard dimensions 40 x 10 x 1 mm. The measurement mode is selected as the tensile mode, and the frequency mode is chosen as sinusoidal oscillation, providing a temperature sweep from $25$ to $80^{\circ}C$ with a heating rate of $2^{\circ}C/min$. The resulting graph depicts the damping parameter against temperature for four different frequency ranges (1, 2, 5, and 10 Hz), as illustrated in Fig. S7a The graph for PLA represents the measure of material damping/energy dissipation concerning temperature, with the peak indicating the glass transition temperature of the material. Fig. S7b represents the variation of storage modulus with temperature. The graph property at low temperatures corresponds to a high-frequency range, as depicted in Fig. S7. The damping parameter remains uniform up to $50^{\circ}C$, signifying that for high frequencies, the damping of the material is also uniform. Therefore, the occurrence of the bandgap in the lattice structure at high-frequency ranges is attributed to the lattice geometry rather than the material itself.
C Finite Element Analysis (FEA) Simulations
Figure S8: Comparison of properties for fractal (a) Bandgap range comparison for different order (n=0,1,2) of fractal structure (b) Force-deflection relation for given fractals.
Figure S9: Comparison of (a) antysymmetric, and (b) symmetric structure and their respective FRF plots.
FEA simulation was performed using Ansys software, focusing on modal and harmonic response modules. The primary preparation of the model involved material selection (PLA) and mesh definition. In the structural modeling, tetrahedron mesh size of 2 mm was chosen based on the geometry’s size and complexity. The mode superposition method links the modal analysis results with the harmonic analysis input. Fixed edge boundary conditions were applied to the face receiving the vibration waves. In the modal analysis, 100 modes were extracted to obtain a broader spectrum of input frequencies for the harmonic response analysis, leading to more accurate results. A dynamic force of amplitude 10 N and frequency varying from 0 to 10 KHz was applied to the same face for vibrational excitation and a damping factor of 0.001 was incorporated for practical considerations.
C.1 Effect of fractal order and lattice symmetry on bandgap
Fractal structures represent self-similar properties, meaning their complex geometry repeats at progressively smaller scales. These self-similar features have no inherent limit and become increasingly intricate with higher fractal orders. Figure S8a illustrates the comparison between different fractal orders and the resulting effect on the bandgap frequency range. Figure S8 shows how the base geometry transitions from a simple honeycomb structure (order 0) to a first order structure with three repeating honeycombs to a second order structure [69, 70]. As the fractal order increases, the bandgap shifts towards higher frequencies, and multiple bandgaps also appear. In addition, the stiffness of the structure also increases with increasing fractal order.
Figure S9 represents the effect of symmetric and antisymmetric structures on wave transmission. To investigate this effect, we have created two structures, one with antisymmetry in both longitudinal and transverse directions, and the other with symmetry in both directions. It clearly demonstrates that an antisymmetric structure is more effective for vibration isolation applications. Apart from antisymmetry, by increasing the waviness of the structure, the bandgap also increases.
The non-dimensional parameter, Gap-mid ratio ($BG$ ), serves as an expression for the width of the bandgap and the center frequency. It provides a measure of the relative size of the bandgap with respect to the band structure. A higher gap-mid-ratio indicates a larger and potentially more useful bandgap.
$BG\%\ =2\frac{(f_{u}-f_{l})}{(f_{u}+f_{l})}\%$ (2)
where $f_{l}$ and $f_{u}$ are the lower and upper-frequency limit of the band gap.
D Experimental Method
Figure S10: (a) The experimental setup, accompanied by the Laser Doppler Vibrometer (LDV), is depicted in the schematic diagram. (b) 3D laser head with the controller is featured. (c) Power amplifier (d) Electro-dynamic shaker which is employed in setup (e)-(f) Excitation is provided from the base through an electrodynamic shaker, while response is obtained by Laser Doppler Vibrometer focused on the top surface of the work-piece.
The study utilizes a non-contact technique to measure vibration to assess the response of different latticed metastructures. Nine 3D printed samples, each showing unique metastructure, are subjected to dynamic testing using a 3D Laser Doppler vibrometer (LDV) from Polytec (Details of the experimental setup presented in Fig. S10). 3D printed structures are made of PLA (the detailed properties of the material are given in Table S3). The properties and composition of the hexagonal shape inserts made of mild steel are shown in Tables S3 and S4. These samples are attached to an Electrodynamic shaker system in cantilever mode. The laser head of LDV captures surface velocity by analyzing the phase shift between the incident and reflected laser beams. Data acquisition and signal processing are executed through the NI-DAQ system. The base excitation technique is utilized to determine transmissibility. In this method, the LDS shaker generates 3200 Fast Fourier Transform (FFT) lines, and the pseudo-random excitation with a frequency range of 0 to 10 KHz. Displacements of the top and bottom surfaces of the samples are recorded to calculate transmissibility after post-processing of the data. A retro-reflective tape is also attached to the surface where data has been recorded to reduce the scattering of the laser beam from the surface and three closely spaced points were taken on both the top and bottom surface for averaging of data to reduce error.
D.1 Damping characterization of interlaced lattice metastructures using half-power bandwidth method
Figure S11: A comparison of damping parameters obtained from dynamic testing performed under base excitation in LDV for different lattice structures (LS1-LS9).
Table S5: The damping value $\zeta$ obtain from dynamic testing performed in LDV.
Lattice No. -3dB Points (Hz) Damping ratio ($\zeta$) Quality factor (Q)
LS1 816.7 828.8 0.007386 67.69
LS2 896.9 903.4 0.003642 137.30
LS3 1083.4 1093.1 0.004481 111.59
LS4 1062.3 1075.0 0.005951 84.02
LS5 785.5 797.0 0.00732 68.30
LS6 855.2 865.0 0.005654 88.43
LS7 876.8 896.4 0.005434 92.01
LS8 738.0 744.5 0.004415 113.25
LS9 1191.8 1196.2 0.001867 267.82
Damping in mechanical systems holds significant importance and is represented in numerous formats. The most common terms are the amplification factor, also known as the quality factor ($Q$), and the viscous damping ratio often expressed as the fraction of critical damping ($\zeta$). The damping parameter was calculated using the half-power bandwidth method based on the damping test results obtained from Laser Doppler Vibrometry (LDV). The test utilized 3200 Fast Fourier Transform (FFT) lines and was driven by a shaker with a unity gain setting for each sample. Data acquisition and signal processing were performed using NI-DAQ.
The highest natural frequency is observed for honeycomb-auxetic (LS9) metastructure (1193.75 Hz) and the lowest natural frequency is observed for ring interlaced (LS5) metastructure (787.5 Hz). Using the half-power bandwidth method, the damping for each parameter was calculated as shown in Fig. S11. The highest and lowest damping observed for honeycomb-honeycomb structure (LS1) and honeycomb auxetic (LS9) metastructure are $\zeta=0.007386$ and $\zeta=0.001867$ respectively (as shown in Table S5).
The formula relates two parameters:
$Q=\frac{1}{2\zeta}$ (3)
The Q value is equivalent to the magnitude of the peak transfer function observed in a single degree of freedom subjected to base excitation at its natural frequency. In this context, we introduce the quality factor or Q factor, which characterizes the system’s bandwidth in relation to the resonant frequency. A low Q factor reduces the significance of the resonant peak, which poses challenges in measuring changes in resonant frequency. Alternatively, the Q factor is defined as the ratio of the energy stored in the system to the energy dissipated per cycle. A higher Q factor (L9) leads to a lower rate of energy loss, resulting in a sharper peak of resonant frequency.
$2\zeta=\frac{\omega_{2}-\omega_{1}}{\omega_{n}}$ (4)
E AI-driven Inverse Design of Metastructures
Figure S12: Metastructure of primary geometry as honeycomb and secondary geometry has honeycomb, auxetic, star, circle, and hierarchical honeycomb used to design different combinations of metastructures for analysis purpose.
In Figure S12 we have selected the honeycomb-based metastructure as the primary structure and six different structures as the secondary structure interlaced inside the honeycomb structure. With the help of different combinations of secondary structures with primary metastructure, we have generated 720 different metastructures for analysis purposes, out of which one metastructures we can see as a resultant geometry. The analysis data has been taken for learning of forward and reverse models.
E.1 Forward analysis model
Figure S13: Main block diagram describing the forward analysis model. LN: Layer Normalization, GA: Geometric Attention.
The Forward analysis block learns the geometrical patterns and their effect on the transmissibility of the metastructure. The model is designed to capture geometrical features at multiple levels of hierarchy in terms of the sizes of patterns. The block diagram of the forward model is shown in Fig. S13. The box shown in the red dotted line is the main geometric feature extractor which is repeated N-times to make the model deeper. The main constituent of the geometric feature extractor is geometric attention (GA). Layer normalization (LN) before the GA and after the summing block is provided to stabilize the training process by reducing internal covariate shifts. Therefore, it helps the model to converge faster. Dropout after the GA block is included for the weight regularization.
Geometric Attention (GA): Fig. S14(a) shows the schematic diagram of the GA block. The GA block contains the FEM-inspired spatial attention (FSA) block to capture the dependencies on adjacent connections in the input geometrical structure. To capture the interlaced complexity of the metastructure, multiple parallel blocks of FSA with convolutional modules with different kernel sizes are employed. The later part of the FSA is the spatial attention module [61]. The functioning of the spatial block to transform the features for geometry learning is described by Fig S14(b)
Figure S14: Block diagram of the geometric attention (GA). (a) Detailed block diagram of GA and (b) the FEM-inspired spatial attention (FSA).
Regression Head: The last block of the forwards analysis model is the regression head, the detailed block diagram is represented in Fig. S15. The purpose of this unit is to process the geometric features from the geometric feature extractor and produce the transmissibility as output. The regression head first transforms the geometric features to represent with reduced dimension. Then, it is flattened and processed by two fully connected (FC) layers with Parametric Rectified Linear Units (PReLU) activation. The second FC is provided with material properties along with geometric features. These material properties are modulus of elasticity $(E)$, Density $(\rho)$, Poisson’s ratio $(\nu)$, and material damping $(d)$. The output of the last-second FC layer is applied to layer normalization (LN) to stabilize the learning. Finally, the last FC layer with linear activation converts the features to the transmissibility having dimension (1000, 1)10001(1000,\;1)( 1000 , 1 ).
Figure S15: Detailed block diagram of the regression head of forward analysis model.
E.2 Inverse design model
Figure S16 shows the main block diagram of the inverse design network. It mainly consists of spectrum self-attention to convert transmissibility to latent space, down sampling network, multiscale residual network to produce interlaced complex structure, and finally upsampling through transpose convolution. Reflection padding of 3 pixels on each side is provided to increase the size of the latent space with reflected boundary. Details of the spectrum to latent space conversion using self-attention and the multiscale residual network are shown in Fig. S17.
Figure S16: Main block diagram describing the implementation of the inverse design model.
Spectrum Self-Attention: The transmissibility spectrum is the frequency response that represents the geometrical properties as well as the material properties of the metastructure. To represent the input transmissibility into 2D latent space, the transmissibility spectrum is given to a multi-head self-attention mechanism followed by a multilayer perceptron (MLP) as shown in Fig. S17(a). Since the transmissibility is the amplitude versus frequency spectrum, a position encoding must be added to represent the absolute position of the amplitude against the frequency in the input sequence. A position encoding of the dimension equal to the system model dimension is defined using alternate sine and cosine functions to represent the even and the odd positions respectively. For more details, readers can refer to “Attention Is All You Need” [71]
Since the objective is to transform the transmissibility spectrum into high-dimensional latent space, first the spectrum is projected to higher dimensional space (the selected embedding dimension is $(1,2048)$ using a fully connected (FC) linear layer. Then, the position term is embedded with position encoding as described above. The spectrum with position encoded is now passed through a multi-head self-attention with the head dimension equal to the embedding dimension. Now the attention output is non-linearly transformed to a further higher dimension to make it suitable for reshaping to the latent space dimension of $(3,128,256)$. The ReLU activation function is for non-linear transformation in the MLP block.
Figure S17: Detailed block diagram of (a) spectrum self-attention (b) multiscale residual network.
Multiscale Residual Network: The reconstruction of metastructure from the latent space produced by the multi-head self-attention module needs transformation of the features of the latent space with multiple scale factor to produce the pattern of multiple hierarchies. Therefore, a residual network with multi-scale kernel sizes of the convolutional module is implemented. The main workflow includes passing the input into different convolutional paths with varying kernel sizes. The selected kernel sizes for our implementation are $(1\times 1)$, $(3\times 3)$, and $(5\times 5)$. This selection is based on the requirement to capture the metastructure at different hierarchies. Each convolutional path consists of two convolutional layers in series with ReLU activation. Instance normalization is provided between the two convolutional modules. The convolutional path with kernel size $(1\times 1)$ helps the model to capture the fine-grained features to construct the tertiary complexity of the metastructure. Similarly, the convolutional path with kernel size $(3\times 3)$ captures the spatial features to construct the secondary geometrical structures. The third path with kernel size $(5\times 5)$ helps the model to extract the broader contextual information for reconstructing the main blocks on the metastructure geometry. Instance normalization is applied to normalize the feature maps across the spatial locations only, improving the stability and efficiency of the training process.
Since the outputs of each path have to be finally concatenated to merge the transformed feature, the dimensionality of the final outcome of each path has to be the same. Therefore, reflection paddings with suitable padding sizes are used in each path before the convolutional layer to adjust the dimensionality with reflected padding. The path with the convolutional kernel of $(1\times 1)$ does not alter the dimensionality of the input space, no padding is provided in this path. To have the same output dimensionality in the convolutional path with kernel sizes $(3\times 3)$ and $(5\times 5)$, the two reflection padding have sizes of $1$ and $2$, respectively.
The residual connections, also known as skip connections (additive connections) are provided after two convolutional operations. This preserves the original input signal, aiding gradient flow and addressing vanishing gradient issues. Since the residual connections add the original features directly to the transformed features to produce the required gradients in the image pattern, again instance normalization is provided to provide stability in the training process.
Finally, the outputs from different kernel paths are concatenated to combine multiscale features into a unified representation. The concatenated feature representation is now projected back to the same spatial dimensionality by using a convolution-based projection layer to represent the final outcome with the same dimensionality. Thus, the resultant output retains both the patterns obtained by multiscale transformation and the residual characteristics of the original input. Concatenating outputs from multiple paths may lead to an increase in feature dimensions and the need for an additional projection layer, thus adding the higher complexity of the model. however, it helps greatly to capture multiple geometrical complexities in the metastructures.
E.3 Loss function
The three loss functions used to define the total loss for the inverse model training are described below.
Mean Square Error (MSE): The mean square error (MSE) can be calculated by taking the difference between the actual and predicted parameters from the following expression
$\mathcal{L}_{1}=\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}-\bar{Y_{i}}\right)^{2}$ (5)
Structural Similarity Index Measure: The structural similarity index measure is a yardstick for comparing two images [64]. SSIM assesses the similarity between images based on structural information available within the image through the correlation of pixel intensities. The SSIM is calculated by considering three components called luminance (l), contrast (c), and structure (s) through the following formula
$\text{SSIM}=\frac{\left(2\mu_{x}\mu_{y}+C_{1}\right)\left(2\sigma_{xy}+C_{2}\right)}{{\left(\mu_{x}^{2}+\mu_{y}^{2}+C_{1}\right)}{\left(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2}\right)}}$ (6)
where $x$ and $y$ refer to images being compared. $\mu_{x}$ and $\mu_{y}$ are mean intensities of images. $\sigma_{xy}$ is covariance of both images. $C_{1}$ and $C_{2}$ are constants included to avoid instability. The higher the value of SSIM, the better the match between the two images. Therefore, the loss to minimize for better learning can be defined as
$\mathcal{L}_{2}=1\;-\;\text{SSIM}$ (7)
The loss term $\mathcal{L}_{2}$ is included in the total loss with a penalty factor of $\beta_{1}$, a positive fractional number.
Maximum Mean Discrepancy (MMD): Maximum mean discrepancy MMD is a statistical method used to calculate the difference between two probability distributions. MMD basically gives the measure of the non-parametric distance between the mean of two distributions on reproducing kernel Hilbert space (RKHS) [72]. Hence, MMD compares the distributions of two sets of data by mapping them into a higher-dimensional feature space using a kernel function, and then finally, it measures the distance between their means in the new feature space. If the means are equal, it implies that the distributions are similar; if the means are far away, then the distributions differ from each other. For a universal RKHS, MMD asymptotically tends to zero if the probability distribution of the two data patterns is the same.
If $x_{i}^{s}\in\textbf{X}^{s}$ and $x_{j}^{t}\in\textbf{X}^{t}$ are $i^{th}$ source data and the $j^{th}$ target data in the source space $\chi^{s}$ and the target space $\chi^{t}$ respectively, then the MMD of the two distribution in RKHS is defined from the below expression
$\mathcal{L}_{3}(\textbf{X}^{s},\textbf{X}^{t})=\left\|{\frac{1}{{n^{s}}}\sum\limits_{i=1}^{{n^{s}}}{f\left({{x_{i}^{s}}}\right)}-\frac{1}{{n^{t}}}\sum\limits_{j=1}^{{n^{t}}}{f\left({{x_{j}^{t}}}\right)}}\right\|_{\mathcal{H}}$ (8)
where, $f(.)$ denotes the kernel function $f:X,X\to{\mathcal{H}}$, ${\mathcal{H}}$ be the universal RKHS, $n^{s}$ and $n^{t}$ be the number of samples in the source and the target datasets $\textbf{X}^{s}$ and $\textbf{X}^{t}$ respectively.
For our case, loss due to MMD is calculated between the original metastructure and the predicted metastructure. It is included in the total loss with a multiplication factor of $\beta_{2}$, which is a positive fractional number. The inclusion of MMD in the total loss term forces the algorithm to minimize the discrepancy
F Image Enhancement for Noise Removal
The metastructure generated by the inverse design model contains lots of colored noises. These noises must be removed for a clear display of the metastructure geometry. The following steps are followed to remove the noises from the generated metastructure.
  1. i)
    Convert to Grayscale: Convert the input image $I(x,y)$ to a grayscale image $Gr(x,y)$ using:
    $Gr(x,y)=0.299R(x,y)+0.587G(x,y)+0.114B(x,y)$ (9)
    where, $Gr(x,y)$ is the spatial dimension of the grayscale image. Similarly, $R(x,y)$, $G(x,y)$, and $B(x,y)$ represent the two-dimensional space of reg, green, and blue channels of the image, respectively.
  2. ii)
    Apply mask to isolate the structure: Create a binary mask $M_{\text{gray}}(x,y)$ using a gray color threshold value:
    $M_{\text{gray}}(x,y)=\begin{cases}255&\text{if }Gr(x,y)<T_{\text{gray}}\\ 0&\text{otherwise}\end{cases}$ (10)
    Apply the binary mask defined above to extract the gray-colored structural regions of the whole image
    $S_{\text{gray}}(x,y)=G(x,y)\frac{M_{\text{gray}}(x,y)}{255}$ (11)
  3. iii)
    Detect black region: Define the lower and upper thresholds for black regions as follows:
[…]
    Also, create a binary mask $M_{\text{black}}(x,y)$ for black regions using:
    $M_{\text{black}}(x,y)=\begin{cases}255&\text{if }Tb_{\text{lower}}\leq I(x,y)\leq Tb_{\text{upper}}\\ 0&\text{otherwise}\end{cases}$ (13)
  4. iv)
    Apply black mask: Apply $M_{\text{black}}(x,y)$ to preserve non-black regions:
    $R_{\text{non-black}}(x,y)=I(x,y)\frac{M_{\text{black}}(x,y)}{255}$ (14)
  5. v)
    Combine the results: Combine $R_{\text{non-black}}(x,y)$ and $S_{\text{gray}}(x,y)$ to produce the final image:
    $F(x,y)={w}_{1}R_{\text{non-black}}(x,y)+{w}_{2}S_{\text{gray}}(x,y)$ (15)
    Here, $w_{1}=1$ and $w_{2}=1$ are weight parameters to combine the masked image data.
  6. vi)
    Normalization: Finally, normalize $F(x,y)$ to the 8-bit range and display as image:
    $F(x,y)=255\left(\frac{F(x,y)-F_{\min}}{F_{\max}-F_{\min}}\right)$ (16)
Appendix 2 Supplementary Discussion
Additive manufacturing of material provides an edge to print some intricate and complex structures precisely for experimental validation, which is difficult to do with traditional machining processes[73, 74]. Within the regular lattice, the lattice metamaterial has a complex structural geometry designed to effectively control and manipulate vibration waves that pass through it, influencing the material’s overall properties. The interlaced structure gives us an edge by providing more parameters to generate a variety of structures. By changing geometric properties, we can make these structures tuneable. Additionally, the application of local resonators allows us to tune the bandwidth for more specific regions of interest. For effective analysis of the materials, the DMA test of PLA and the XRF test of the mild steel sample has been done. For further analysis of the interlaced structure, a self-similar interlaced structure which is also called a fractal, similar FEM analysis has been done. These properties are essential to designing vibration isolators for sensitive structures.
A total of 720 data samples with honeycomb-based interlaced metastructure and corresponding transmissibility spectrum have been prepared to develop an AI-based inverse design of metastructure. This development is accomplished in two phases of training of deep learning models. In the first phase, the images of metastructure geometry were used to train a forward analysis model to predict the corresponding transmissibility spectrum. The forward analysis model works based on geometrical feature extraction using the multi-head GA module. The multi-head GA has been designed to learn the hierarchies of the metastructure geometry in terms of multiple kernel sizes of convolutional-based feature transformation. The second phase of the training involves the development of an inverse design deep learning model that uses the transmissibility spectrum to produce the corresponding metastructure. The inverse design model has been proved to be a breakthrough in designing a metastructure for the desired frequency bandgap. Therefore, given the desired frequency bandgap, an artificial transmissibility spectrum can be produced and applied to the inverse design model to generate a suitable metastructure that exhibits the frequency response of the desired bandgap. This also opens doors for exciting applications in areas such as vibration damping, energy absorption, and even the creation of materials with programmable mechanical behavior.

### v3

@@
-Inverse design of potential metastructures inspired from Indian medieval architectural elements
+AI-driven Inverse Design of Band-Tunable Mechanical Metastructures for Tailored Vibration Mitigation
 Abstract
-In this study, we immerse in the intricate world of patterns, examining the structural details of Indian medieval architecture for the discovery of motifs with great application potential from the mechanical metastructure perspective.
-The motifs that specifically engrossed us are derived from the tomb of I’timad-ud-Daula, situated in the city of Agra, close to the Taj Mahal.
-In an exploratory study, we designed nine interlaced metastructures inspired from the tomb’s motifs.
-We fabricated the metastructures using additive manufacturing and studied their vibration characteristics experimentally and numerically.
-We also investigated bandgap modulation with metallic inserts in honeycomb interlaced metastructures.
-The comprehensive study of these metastructure panels reveals their high performance in controlling elastic wave propagation and generating suitable frequency bandgaps, hence having potential applications as waveguides for noise and vibration control.
-Finally, we developed novel AI-based model trained on numerical datasets for inverse design of metastructures with a desired bandgap.
-Keywords: Metamaterials, Metastructures, Interlaced metastructures, Hierarchical Metastructures, Fractals, Honeycomb, Elastic wave control, Artificial Intelligence, Multihead geometric attention, Multiscale residual network, Inverse design.
-Introduction
-Architectural intricate patterns have a long history, with certain examples dating back to the 4th century [1].
-However, these early forms were the outcome of creative heuristic advancements of pattern generation in human civilization.
+On-demand vibration mitigation in a mechanical system needs the suitable design of multiscale metastructures, involving complex unit cells.
+In this study, immersing in the world of patterns and examining the structural details of some interesting motifs are extracted from the mechanical metastructure perspective.
+Nine interlaced metastructures are fabricated using additive manufacturing, and corresponding vibration characteristics are studied experimentally and numerically.
+Further, the band-gap modulation with metallic inserts in the honeycomb interlaced metastructures is also studied.
+AI-driven inverse design of such complex metastructures with a desired vibration mitigation profile can pave the way for addressing engineering challenges in high-precision manufacturing.
+The current inverse design methodologies are limited to designing simple periodic structures based on limited variants of unit cells.
+Therefore, a novel forward analysis model with multi-head FEM-inspired spatial attention (FSA) is proposed to learn the complex geometry of the metastructures and predict corresponding transmissibility.
+Subsequently, a multiscale Gaussian self-attention (MGSA) based inverse design model with Gaussian function for 1D spectrum position encoding is developed to produce a suitable metastructure for the desired vibration transmittance.
+The proposed AI framework demonstrated outstanding performance corresponding to the expected locally resonant bandgaps in a targeted frequency range.
+keywords:
+Metamaterials , Interlaced metastructures , Fractals , Elastic wave control , Geometric feature extraction , Multihead FEM-inspired spatial attention , Multiscale residual network.
+journal: Mechanical Systems and Signal Processing
+[…]
+organization=Department of Mechanical Engineering, Indian Institute of Technology Kanpur,postcode=208016, country=India
+[…]
+organization=Department of Precision and Microsystems Engineering, Delft University of Technology,postcode=2628 CD, country=Netherlands
+[…]
+organization=Department of Mechanical Engineering, Motilal Nehru National Institute of Technology Allahabad,city=Prayagraj, postcode=211004, country=India
+1 Introduction
+Metamaterials are emerging from varied research domains like electromagnetic, mechanical, and acoustic applications having unique tailored properties that are primarily dependent on their structural geometry rather than intrinsic material properties.
 It wasn’t until the late 19th century, when fundamental physics became better understood, that the development of systematic engineering principles made it possible to explore counter-intuitive physical characteristics of some patterns in periodic structures, generally known as metamaterials and metastructures.
-Today, it is often proposed as a panacea for controlling attenuation and wavefront divergence related to elastic and electromagnetic wave propagation and there exists a vast literature of which are based on meticulously designed lattice structures with special geometric patterns [2, 3, 4, 5, 6].
-These engineered lattices may introduce exotic mechanical and electromagnetic properties, manipulating the behavior of electromagnetic waves [7], elastic waves as well as sound waves [8, 9, 10, 11].
-By carefully designing these lattices, scientists can achieve extraordinary effects like negative stiffness [12], negative mass [13], and negative Poisson’s ratio [14].
-The study of lattice geometries in metamaterials opens a pathway to myriad innovative applications ranging from optics to acoustics, wearable devices [15, 16], advanced space structures [17, 18], next-gen automobiles [19, 20], robotic manipulators [21], bionic grippers, [22] and anti-seismic structures [23, 24].
-These metamaterials exhibit characteristics such as high specific stiffness [25, 26, 27, 28], negative Poisson’s ratio [29, 30], negative thermal expansion coefficient [31], negative reflection [32, 33], wave alteration [34], sub-wavelength bandgap [35, 36] and distinctive electro-dynamic phenomena [37, 38].
-The generation of a large range of novel periodic structures at different length scales has ushered in phenomenal developments of metamaterials.
-Figure 1: The Tomb of Itmad-ud-Daula is a Mughal mausoleum in the city of Agra in the Indian state of Uttar Pradesh (27.1929 N, 78.0310 E).
+Today, it is often proposed as a panacea for controlling attenuation and wavefront divergence related to elastic and electromagnetic wave propagation and there exists a vast literature based on meticulously designed lattice structures with special geometric patterns [1, 2, 3, 4, 5, 6].
+Generally, these structures have sub-wavelength features designed in the form of lattices on a smaller scale than the wavelength of external physical stimuli they interact with.
+These engineered lattices may introduce exotic mechanical and electromagnetic properties, manipulating the behavior of electromagnetic waves [7], elastic waves as well as sound waves [8, 9, 10].
+As metastructures have structural-dependent properties, hence by carefully designing these lattices, scientists can achieve customizable responses with extraordinary effects like negative stiffness [11, 12, 13], negative mass [14], negative Poisson’s ratio [15, 16], negative thermal expansion coefficient [17], negative reflection [18, 19] for wave alteration [20], sub-wavelength bandgap [21, 22, 23] and distinctive electro-dynamic phenomena [24, 25, 26].
+The study of lattice geometries in metamaterials opens a pathway to myriad innovative applications ranging from optics to acoustics such as wearable devices [27, 28], advanced space structures [29, 30], next-gen automobiles [31, 32], high precision robotic manipulators [33], bionic grippers, [34], anti-seismic structures [35, 36], and vibration sensing [37].
+Mechanical properties of metastructures designed with simple unit cells such as honeycombs with positive and negative re-entrant angles, chiral, and kagome have been extensively studied through experimental avenues, numerical simulations, and mathematical analyses [38, 39, 40, 41, 42, 43, 44].
+The generation of a large range of novel periodic structures at different length scales has ushered in phenomenal developments of metamaterials as depicted in Fig.
+1.
+Figure 1: Different types of metastructure unit cells can be categorized based on their lattice structures.
+(a) Five popular unit cells are used as primary lattices.
+(b) A complex array of primary lattices can be used to create different possible interesting metastructure unit cells.
+(c) Multi-unit patterns represent the arrangement of unit cells to form complete 2D and 3D metastructures.
+Architectural patterns in metamaterials can also play a pivotal role in manipulating electromagnetic and acoustic waves.
+Intricate architectural patterns have a long history, with certain examples dating back to the 4th century [45].
+These early forms were the outcome of creative heuristic advancements of pattern generation often driven by culture and religion.
+Geometric design and pattern have played a significant role in the spiritual traditions of many religions.
+Considered as the inspiration for the Taj Mahal, the tomb of Itmad-ud-Daula, also known as the ‘Baby Taj Mahal,’ is a stunning example of Indian medieval architecture as shown in Fig.
+2 (a).
+This mausoleum displays a wonderful variety of geometric arabesques.
+Such intricate patterns, popularly known as ‘Jali’ in India, were crafted for aesthetic purposes and also as a functional component in passive cooling and ventilation.
+Some of the patterns are shown in Fig.
+2(b-f) with the corresponding metastructural models shown in Fig.
+2(g).
+The related lattice structures can be classified according to different types of unit cells, considering angle, arrangement, and symmetry (Fig.
+1(a)).
+We investigate various complex/hybrid arrangements of these unit cells (Fig.
+1(b)) and then demonstrate how different arrangements can be used to generate complete metastructures (Fig.
+1(c)).
+These metastructures provide a foundation for further studies evaluating their performance in controlling elastic wave propagation and vibration.
+When we delve into the realm of complex hybrid unit cell-based metastructure it requires advanced manufacturing techniques like 3D printing.
+Recent advancements in precise additive manufacturing techniques at the microscale have further enabled the fabrication of new lattices with complex topology and interlaced geometry, which are challenging to achieve using traditional subtractive manufacturing methods.
+Such metastructures have distinct behaviors when subjected to external stimulus such as mechanical excitation [46, 47, 48, 49, 50].
+Additive manufacturing has the potential to fabricate precisely complex cellular and architectural metamaterials by manipulating multimaterials with controllable properties.
+The design and fabrication of smart material-based programmable mechanical metamaterials with 4D printing technology can change their properties dynamically in response to external stimuli.
+In recent years, there has been a growing interest in developing engineered materials using a range of 3D and 4D additive manufacturing techniques [51, 52].
+Such metastructures with complex multiscale geometry are difficult to design according to a desired frequency response with a targeted bandgap.
+It may require numerous attempts of FEM-based simulations and design changes to achieve the targeted frequency response and, hence, may not be feasible in practice due to high computational costs.
+Therefore, efficient AI-based models are necessary that can predict the effective design of a metastructure with complex geometry and the desired frequency response.
+In recent years, various attempts have been made to develop AI-based inverse design models that can generate complex multiscale mechanical metamaterials having tailored mechanical response and properties [53, 54, 55].
+Deng et.
+al.
+[56] suggested a neural accelerated evolutionary strategy that can obtain a geometry of metamaterials with hinged quadrilaterals for a wide range of target nonlinear mechanical responses.
+Figure 2: (a)The Tomb of Itmad-ud-Daula is a mausoleum in the city of Agra in the Indian state of Uttar Pradesh (27.1929 N, 78.0310 E).
 The architecture of this monument is unique in terms of crafting which uses a combination of pietra-dura and calligraphy-based decorative art techniques.
-It is considered to be the pioneer of the creation of funerary tombs in Mughal-India which was subsequently magnified in the Taj Mahal.
-However, with its marble latticework and exquisite carvings, the tomb is considered as a work of art even more delicate than the Taj Mahal.
 (b to d).
@@
 Here, we have specifically focused on the Jali screens at the centre of the niches all around the tomb.
-They were used extensively during the Mughal reign for elegance and their ability to facilitate the entry of natural light and air inflow.
+They were used extensively for elegance and their ability to facilitate the entry of natural light and air inflow.
 Often it is also used as part of a passive cooling system.
-(g) shows the corresponding structural models (LS1-LS9).
-Figure 2: Different types of metastructure unit cells can be categorized based on their lattice structures.
-(a) Five popular unit cells are used as primary lattices.
-(b) A complex array of primary lattices can be used to create different possible interesting metastructure unit cells.
-(c) Multi-unit patterns represent the arrangement of unit cells to form complete 2D and 3D metastructures.
-Mechanical properties of simple metastructures such as honeycombs with positive and negative re-entrant angles have been extensively studied through experimental avenues, numerical simulations, and mathematical analyses [39, 40, 41, 42, 43, 44].
-In recent years, there has been a growing interest in developing engineered materials using a range of 3D and 4D additive manufacturing techniques [45, 46].
-Recent advancements in precise additive manufacturing techniques at the microscale have further enabled the fabrication of new lattices with complex topology and interlaced geometry, which are challenging to achieve using traditional subtractive manufacturing methods.
-Such metastructures have distinct behaviors when subjected to external stimulus such as mechanical excitation [47, 48, 49, 50, 51, 52, 53].
-Geometric design and pattern have played a significant role in the spiritual traditions of most of the religions.
-For instance, concentric circles and triangles are widely used as sacred patterns in the Hindu and Buddhist ‘Mandala’; while rectangles and squares are preferred in the ‘Feng Shui’ design of Chinese spiritual tradition.
-The construction of medieval European cathedrals was often based on the ‘Latin Cross’.
-Fractal geometries were the most popular design used in the construction of various Hindu temples.
-The pinnacle of embedding geometric motifs is found in ‘Islamic Geometric Patterns’ which involve more complicated shapes such as hexagons, octagons, decagons, and multi-point stars [54].
-One of the interesting applications of such motifs in tessellated form is called ‘Mashrabiya’ which is known to be originated in Baghdad during the 12th-century Abbasid period.
-Its interlaced arrangements, which emerged from simplicity and unity of natural as well as geometric motifs, exhibit some fundamental relationships that become in-turn hierarchical or fractal in complex architectures.
-The architectural applications of ‘Mashrabiya’ had rapidly developed in the medieval Persian and Indian civilizations due to the influence of Islamic religion and mysticism.
-Such patterns, popularly known as ‘Jali’ in India, were crafted not only for aesthetic purposes but also for their utility in terms of transmitting indirect sunlight, capturing fresh air, passive cooling, maintaining communication with privacy, etc.
-(some of the patterns are shown in Fig.
-1(b-f)).
-The first instance of developing such patterns in India was found in the Alai Darwaza in the Qutb Minar complex of Old Delhi.
-The tomb of Itimaduddaula, a magnificent testament to the Mughal style enriches Agra’s vibrant tapestry.
-Considered as the inspiration for the Taj Mahal, the tomb also known as the ‘Baby Taj Mahal’, is a stunning example of Mughal architecture (Fig.
-1 (a)).
-This mausoleum displays a wonderful variety of geometric arabesques.
-The strong influence of the Persian style coupled with numerous Indian motifs is visible in this structure.
-The related lattice structures can be classified based on different unit cell types, considering angle, arrangement, and symmetry (Figure 2(a)).
-We investigate various interlacing arrangements of these unit cells (Figure 2(b)) and demonstrate how different arrangements can be used to generate complete metastructures (Figure 2(c)).
-These metastructures provide a foundation for further studies evaluating their performance in controlling elastic wave propagation and vibration.
-Further, we have developed an AI-based forward analysis model to predict the transmissibility of a metastructure.
-Also, an AI-based inverse design of metastructure is developed to design new lattices for the desired bandgap creating end-less possibilities of generating suitable patterns focused to an application [55, 56, 57, 58, 59, 60].
+(g) Shows the corresponding metastructure models (LS1-LS9) inspired from these architectural elements.
+In another similar work, Ha et.
+al.
+[57] developed a rapid inverse design method using generative machine learning to generate metamaterials having uniaxial compressive stress-strain response.
+Recently, Bastek et.
+al.
+[58] proposed video-denoising diffusion models that can produce materials with tailored complex target performance.
+Their AI-based model, having spatial and temporal attention, enables the inverse design of metamaterial with a highly non-linear stress-strain response that closely matches with the finite element simulations.
+These research works were dedicated to the inverse design of metamaterials with tailored nonlinear stress-strain properties.
+Similarly, various researchers have also investigated the AI-based inverse design of metamaterials and metasurfaces like non-periodic 3D architectures having predictable anisotropic properties[59], low-index-contrast structures on a silicon photonics platform [60] and antenna surface design [61].
+All these studies have generally focused on the inverse design of unit cells of metamaterials or periodic surfaces for a tailored mechanical or electrical response.
+The inverse design of mechanical metastructures having complex interlaced geometry for tailored frequency response in a defined attenuation bandwidth still remains a challenge due to hierarchical multiscale geometry.
+Therefore, through this study, we propose a FEM-inspired geometric attention mechanism that can learn the geometry having multiple scales in a metastructure and produce the corresponding frequency response.
+Further, with the help of the forward model, an inverse design model is developed that can generate a suitable metastructure for the desired vibration response.
+As mechanical metamaterials have structure-dependent properties hence, their inverse design can revolutionize various application domains with customizable responses by tuning mechanical properties such as stiffness deformation, flexibility, and vibration response.
+Hence AI-based inverse design of mechanical metamaterials can enable advanced functionalities in prosthetics, exoskeletons, and bone implant designs, dynamic and adaptive structure design in robotics, impact and shock energy absorption lightweight metastructures, and soft robotics [62].
+2 Numerical Simulations
+Finite element analysis-based simulation has been carried out using Ansys, employing the Ansys Parametric Design Language (APDL) solver for all 720 structures, and the host material is chosen as PLA (detailed properties given in Supplementary Table S3).
+The entire structure was discretized by using tetrahedron elements with a 2 mm element size, to ensure convergence for interlaced geometries.
+The smaller edges of the base structure were fixed, while the longer edges remained free.
+A dynamic force of amplitude equivalent to 10 N was applied to the base, and the response was calculated on the opposite phase.
+Modal analysis was performed to identify all modes up to 10 KHz frequency, followed by harmonic analysis to obtain the transmissibility.
+The mode superposition method was used to link the modal analysis results with the harmonic analysis input.
+Frequency response functions (FRFs) were generated over the frequency range of 0 to 10 KHz with a 10 Hz frequency resolution; this step size is preferable to accurately capture the response.
+A structural damping ratio of 0.001 was incorporated to account for the viscoelastic effect of PLA substrate.
+2.1 Dynamic characterization and determination of the Bandgap
 Figure 3: Schematic diagram illustrates latticed mechanical metastructures featuring different interlaced lattice structures (LS1-LS9).
 Dynamic analysis of the metastructures is performed and mode shapes of lattices at a frequency corresponding to minimum force transmissibility for in-plane excitation are shown.
-The attenuations corresponding to each geometry are obtained by using the transmissibility plot for each structure.
+In mode shapes, red indicates maximum deformation and blue indicates minimum deformation.
+The attenuation corresponding to each geometry is obtained by using the transmissibility plot for each structure.
 We have highlighted distinct attenuation zones within the transmissibility response of metastructures through green-shaded patches.
+Selected latticed metastructures as shown in Fig. 3, generated based on architectural elements (described in Fig. 2) have an overall size of 204 mm x 58 mm and a thickness of 5 mm.
 Figure 4: Conducting a comparative analysis, this study investigates the impact of scaling (x, 1.5x, 2x) of the unit cell on the bandgap, with the resulting BG ratio illustrated in a bar diagram.
@@
 Subunit cells are used as an interlaced component for the primary structure from which each configuration is constructed.
-The patterns of latticed mechanical metastructures are 3D Printed using PLA having different primary lattice structures namely (a) Honeycomb-honeycomb interlaced (LS1) (b) Honeycomb interlaced with honeycomb pattern (LS2) (c) Honeycomb interlaced with a star (LS3) (d) Chiral interlaced structure (LS4) (e) Ring interlaced structure (LS5) (f) Fishtail structure (LS6) (g) Star-honeycomb structure (LS7) (h) Honeycomb-auxetic structure (LS8) (i) Honeycomb-auxetic interlaced structure (LS9).
-Figure 5: Experimental validation of computational result with variations in mass inserts pattern is shown, and the corresponding bandgaps are marked by green patches.
-(a) Simple interlaced honeycomb structure (LS1) (b) Diamond pattern with a broken symmetry at the middle (c) Cross pattern (d) Forward arrow pattern (e) Diamond pattern (f) Top and bottom row are filled middle row is empty (g) Top and bottom rows are filled with mass inserts while the middle row has two-gap arrangement with no inserts (h) Upward pointing arrow pattern (i) Top and bottom row inserts are arranged alternatively while middle row has two-gap arrangement with no inserts.
-Results
-Dynamic characterization and determination of the Bandgap
-Selected latticed metastructures (represented in Fig.
-3), generated based on architectural elements (described in Fig.
-2) have an overall size of 204 mm x 58 mm and a thickness of 5 mm.
+The patterns of latticed mechanical metastructures are 3D Printed using PLA having different primary lattice structures namely (a) Honeycomb-honeycomb interlaced (LS1) (b) Honeycomb interlaced with a honeycomb pattern (LS2) (c) Honeycomb interlaced with a star (LS3) (d) Chiral interlaced structure (LS4) (e) Ring interlaced structure (LS5) (f) Fishtail structure (LS6) (g) Star-honeycomb structure (LS7) (h) Honeycomb-auxetic structure (LS8) (i) Honeycomb-auxetic interlaced structure (LS9).
 The size was carefully chosen keeping in mind the manufacturing constraints as well as the requirement of minimum number of periodic units (greater than 7 units) in the tessellated structure to express the metamaterial behavior [14].
@@
 3.
-In the middle column (Fig.
-3), we have shown the related mode shapes obtained using Ansys corresponding to minimum force transmissibility within the attenuation zone.
+We have shown the related mode shapes obtained using Ansys corresponding to minimum force transmissibility within the attenuation zone.
 Different structures have different damping ratios depending on the lattice geometry.
-The damping ratio is calculated by the half-power bandwidth method (highest for LS1 and the lowest for LS9), which is shown in the supplementary Fig.
-S11.
+The damping ratio provided is calculated by the half-power bandwidth method (highest for LS1 and the lowest for LS9), which is in the supplementary Fig.
+S10.
 Latticed metastructures with different sub-lattice structures are shown in Fig.
@@
 After creating all the metastructures of specified dimensions, we scaled them up by 1.5 and 2 respectively to capture the scale effect of unit cells in the wave transmission pattern.
-The scaling effect is illustrated with the help of the Band Gap (BG) ratio in the bar diagram (Fig.
-4).
-Performance measurement of lattice with resonators
+The scaling effect as shown in Fig.
+4 is illustrated with the help of the Band Gap (BG) ratio in the bar diagram.
+2.2 Bandgap characteristics of honeycomb-based interlaced structures
+Figure 5: (a) A comparison of bandgap at different frequency levels of different structures interlaced in the honeycomb structure.
+(b), (d) The comparison of static performance is obtained through analytical computations of different structures.
+The displacement control mechanism is set as 5 mm compression of the initial height of samples (c) Comparison of bandgap in different order (n=0, 1, 2) of the fractal structure.
+Comparative analysis of the effect of different interlaced structures on bandgap shift (Fig.
+5(a)) is carried out in various frequency ranges.
+The band gap occurs at a frequency band comparatively higher than that of the simple honeycomb structure.
+Notably, for the LS3 structure bifurcation in bandgap occurs, and for the last structure (LS9) multiple bandgaps occur both for the higher and lower frequency ranges.
+Furthermore, we have compared the bandgap range of various interlaced honeycomb unit cells (Fig.
+5(b)) along with the force-deflection relationships.
+The force-deflection relationships are linear for all the patterns except for LS9.
+Figure 5(c) illustrates the fractal structure of different order (n=0, 1, 2) and their corresponding bandgap shifts.
+In different structures, as the fractal order increases, the bandgap shifts to higher frequency ranges, and for further increase in fractal order multiple bandgaps occur.
+Furthermore, on the right side we have represented the force-deflection relationship for the fractal structure (see Fig.
+5(d)).
+With the increase in the fractal order, the structure becomes stiffer, with 196.32% increase from zeroth to first-order fractal and 33.61% increase from first to second-order fractal.
+From this study, we can conclude that with increasing the structural complexity inside the structure, the bandgap starts to shift towards the higher frequency range, and for higher-order fractals, it also bifurcates which is verified by taking different fractal structures inside the honeycomb and explained in the Supplementary Fig.
+S8.
+3 Experimental testing
+Figure 6: (a) The experimental setup, accompanied by the Laser Doppler Vibrometer (LDV), is depicted in the schematic diagram.
+(b) 3D laser head with a controller is featured.
+(c) Power amplifier (d) Electro-dynamic shaker which is employed in setup (e)-(f) Excitation is provided from the base through an electrodynamic shaker, while the response is obtained by Laser Doppler Vibrometer focused on the top surface of the work-piece.
+Figure 7: Experimental validation of computational result with variations in mass inserts pattern is shown, and the corresponding bandgaps are marked by green patches.
+(a) Simple interlaced honeycomb structure (LS1) (b) Diamond pattern with a broken symmetry at the middle (c) Cross pattern (d) Forward arrow pattern (e) Diamond pattern (f) Top and bottom row are filled middle row is empty (g) Top and bottom rows are filled with mass inserts while the middle row has two-gap arrangement with no inserts (h) Upward pointing arrow pattern (i) Top and bottom row inserts are arranged alternatively while middle row has two-gap arrangement with no inserts.
+The study utilizes a non-contact technique to measure vibration to assess the response of different latticed metastructures.
+3D printed samples, each showing unique metastructure, are subjected to dynamic testing using a 3D Laser Doppler Vibrometer (LDV) from Polytec as shown in Fig.
+6.
+3D printed structures are made of PLA the detailed properties of the material and elemental composition of the hexagonal shape inserts made of mild steel are shown in supplementary Tables S3 and S4.
+These samples are attached to an Electrodynamic shaker system in cantilever mode.
+The laser head of LDV captures surface velocity by analyzing the phase shift between the incident and reflected laser beams.
+Data acquisition and signal processing are executed through the NI-DAQ system.
+The base excitation technique is utilized to determine transmissibility.
+In this method, the LDS shaker generates 3200 Fast Fourier Transform (FFT) lines, and the pseudo-random excitation with a frequency range of 0 to 10 KHz.
+Displacements of the top and bottom surfaces of the samples are recorded to calculate transmissibility after post-processing of the data.
+A retro-reflective tape is also attached to the surface where data has been recorded to reduce the scattering of the laser beam from the surface and three closely spaced points were taken on both the top and bottom surface for averaging of data to reduce error.
+3.1 Performance measurement of lattice with resonators
 In figure 7, we have selected a primary honeycomb lattice structure interlaced with a single honeycomb unit (LS1) at the centre.
@@
 The presence of inserts significantly influences the wave transmission patterns.
-Furthermore, the strategic placement of the inserts are considered in such a manner that the inserts themselves make a pattern inside the lattice structure.
+Furthermore, the strategic placement of the inserts is considered in such a manner that the inserts themselves make a pattern inside the lattice structure.
 By changing their location, we can control the bandgap for planar excitation in 1D.
 In Figure 7(a), we present the effect of random vibration through the simple lattice, while in Fig.
-5(b) we explore the effect of wave transmission in broken symmetry of diamond pattern where the middle row of the lattice is fully filled while the top and bottom row is filled with metallic inserts with a definite pattern.
+7(b) we explore the effect of wave transmission in broken symmetry of diamond pattern where the middle row of the lattice is fully filled while the top and bottom row is filled with metallic inserts with a definite pattern.
 In Fig.
-5(c), the effect of wave transmission in the cross-shape pattern is shown where the inserts are arranged alternatively in the middle row of lattice and the top and bottom row is completely filled.
+7(c), the effect of wave transmission in the cross-shape pattern is shown where the inserts are arranged alternatively in the middle row of the lattice and the top and bottom rows are completely filled.
 In Fig.
-5(d) lattices are filled in such a manner that it forms an arrow indicating towards the right direction.
+7(d) lattices are filled in such a manner that it forms an arrow indicating the right direction.
 Figure 7(e), represents the diamond pattern where the middle row is fully filled; however, the top and bottom rows are alternatively filled.
 In Fig.
-5(f), the top and bottom rows are fully filled and the middle row is fully empty.
+7(f), the top and bottom rows are fully filled and the middle row is fully empty.
 In Fig.
-5(g), the top and bottom rows are fully filled; however, the middle row is filled such that there is a spacing of two units between the consecutive inserts.
+7(g), the top and bottom rows are fully filled; however, the middle row is filled such that there is a spacing of two units between the consecutive inserts.
 In Fig.
-5(h), the top row is alternatively filled starting with a blank cell while the bottom row is also alternatively filled starting with a filled cell, and the middle row is fully filled except for the first and the last cell representing upward pointing arrow.
+7(h), the top row is alternatively filled starting with a blank cell while the bottom row is also alternatively filled starting with a filled cell, and the middle row is fully filled except for the first and the last cell representing upward pointing arrow.
 In Fig.
-5(i), the top and bottom rows are alternatively filled while the middle row is filled with two-unit spacing.
+7(i), the top and bottom rows are alternatively filled while the middle row is filled with two-unit spacing.
 Through the study of dynamic response corresponding to harmonic excitation for all these combinations, we aim to find out which pattern location is more sensitive and affects more at a specific frequency range of wave transmission.
@@
 Further, the experiments corresponding to each iteration are carried out to validate our simulation results by using a Laser Doppler Vibrometer (LDV).
-Random vibration up to 10 KHz is performed using LDS shaker with the gain set as unity.
-Detailed experimental setup can be seen in Supplementary Fig.
-S10.
-As may be observed in Fig.
-5, our experimental results show good agreement with the simulation.
+Random vibration up to 10 KHz is performed using LDS shaker with gain set as unity.
+Detailed experimental setup can be seen in Fig.
+6.
+As observed in Fig.
+7, our experimental results show good agreement with the simulation.
 The printing pattern, and infill density (presented in Supplementary Table S2), also affect the attenuation band.
-Bandgap characteristics of honeycomb based interlaced structures
-Figure 6: (a) A comparison of bandgap at different frequency levels of different structures interlaced in the honeycomb structure.
-(b), (d) The comparison of static performance is obtained through analytical computations of different structures.
-The displacement control mechanism is set as 5 mm compression of initial height of samples (c) Comparison of bandgap in different order (n=0,1,2) of the fractal structure.
-Comparative analysis of the effect of different interlaced structures on bandgap shift (Fig.
-6(a)) is carried out in various frequency ranges.
-The band gap occurs at a frequency band comparatively higher than that of the simple honeycomb structure.
-Notably, for the LS3 structure bifurcation in bandgap occurs, and for the last structure (LS9) multiple bandgaps occur both for the higher and lower frequency ranges.
-Furthermore, we have compared the bandgap range of various interlaced honeycomb unit cells (Fig.
-6b) along with the force-deflection relationships.
-The force-deflection relationships are linear for all the patterns except for LS9.
-Figure 6(c) illustrates the fractal structure of different order (n=0,1,2) and their corresponding bandgap shifts.
-In different structures, as the fractal order increases, the bandgap shifts to higher frequency ranges, and for further increase in fractal order multiple bandgaps occur.
-Furthermore, on the right side we have represented the force-deflection relationship for the fractal structure (see Fig.
-6d).
-With the increase in the fractal order, the structure becomes stiffer, with 196.32% increase from zeroth to first-order fractal and 33.61% increase from first to second-order fractal.
-From this study, we can conclude that with increasing the structural complexity inside the structure, the bandgap starts to shift towards the higher frequency range, and for higher-order fractals, it also bifurcates which is verified by taking different fractal structures inside the honeycomb and explained in the Supplementary Fig.
-S8.
-AI-driven Inverse Design of Metastructures
+4 AI-driven Inverse Design of Metastructures
 With the phenomenal expansion potential of the geometry of metastructures, it is possible to use AI efficiently for both forward analysis in predicting the bandgaps as well as inverse design where for a given bandgap a suitable pattern may be obtained.
-The complete flow diagram of a possible AI-based inverse design of metastructures is depicted in Fig.
-7.
-The entire method comprises two components: (i) the forward analysis model and (ii) the inverse design model.
-The forward analysis model acts as a surrogate model for training the inverse design model.
-It is trained independently for the forward analysis of the metastructures with known transmissibility.
+The complete flow diagram of the proposed AI-based inverse design of metastructures is depicted in Fig.
+8.
+The entire method comprises two components: (i) the FEM-inspired spatial attention (FSA)-based forward analysis model and (ii) the multiscale Gaussian self-attention (MGSA)-based inverse design model.
+The forward analysis model acts as a surrogate model for training the inverse design model in the tandem structures of neural networks.
+The forward analysis model is trained independently using the training samples of metastructure and its transmissibility for the prediction of transmissibility for new metastructures.
 Figure 8: AI-based inverse design of metastructures: (a) Training strategy of forward analysis model to predict the transmissibility for a given metastructure (b) Training strategy of inverse design model to predict the metastructure for given transmissibility spectrum.
-Forward Analysis Model
-The forward analysis model consists of repeated blocks of geometric feature extraction modules capable of learning complex metastructure geometry having primary, secondary, and tertiary complexity.
-Each geometrical feature extraction module consists of a multi-scale geometric attention mechanism having multiple heads of FEM-inspired spatial attention (FSA) blocks.
-The FSA is designed to learn the dependencies of adjacent connections at different scales using different kernel sizes of convolutional modules.
-Each FSA consists of an input convolutional layer for feature transformation followed by a spatial attention block [61, 62].
-The complete block diagram of the forward analysis block, multi-scale GA module, and the FSA block are provided in the supplementary Figures S13 and S14.
-The forward analysis model was implemented with 12 layers of geometric feature extraction blocks and 8 heads of FSA in each block.
-The kernel sizes of the input convolutional layer in 8 FSA heads are selected as $(1,1,1,3,3,3,5,5)$.
-This combination of kernel sizes puts more attention on the smallest structures (tertiary), than the secondary, and the primary structures.
-The output of the geometric features extraction module is processed by a regression head that transforms the geometric features to the transmissibility of dimension $(1000,1)$.
-The detailed block diagram and the descriptions are provided in the supplementary file: Fig.
-S15 and Section E.1.
-Inverse Design Model
-The inverse design model is developed to obtain the metastructure geometry using transmissibility plots for the desired bandgap.
-The detailed block diagram of the inverse design model is provided with the supplementary materials (Figures S16 and S17).
-The key contributions of the inverse design model are the conversion of transmissibility into latent space using a multi-head 1-D self-attention mechanism, followed by constructing the interlaced patterns using a multiscale residual network.
-The 1-D self-attention mechanism captures the sequential dependencies of the transmissibility spectrum, [63] and generates a rectified latent space of shape $(3\times 128\times 256)$.
-The latent space is further extended to a new higher space using reflection padding and then non-linearly down-sampled back to the same dimension using multiple consecutive convolutional layers with ReLU activation.
-Now the transformed latent space is passed through a multiscale residual network having three parallel convolutional paths with three different kernel sizes: $(1\times 1)$, $(3\times 3)$, and $(5\times 5)$ as shown in Fig.
-S17(b).
+Figure 9: Detailed block diagram of the FSA-based forward analysis model: (a) Complete block diagram of the forward model showing the transformation of input metastructure to output transmissibility.
+The orange box shows the geometric attention (GA) model responsible for capturing the geometrical dependency of vibration mitigation.
+Grey boxes show the 8 heads of the FEM-inspired spatial attention (FSA) module.
+(b) Detailed diagram of FSA module.
+The kernel size of the input convolutional block ($f^{\alpha\times\alpha}$) is different for each FSA head for capturing the inter-dependencies of adjacent connections in the metastructure geometry.
+4.1 Forward Analysis Model
+The forward analysis model consists of $N$ blocks of geometric feature extractor (GFE), followed by a regression head to process the output of GFE and the material properties to predict the transmissibility of the metastructure.
+The GFE module is designed on the concept of finite element model (FEM) with the capability of learning the neighborhood dependencies in the complex metastructure geometry having primary, secondary, and tertiary complexity.
+Each GFE module consists of a multi-scale geometric attention (GA) mechanism having multiple heads of FEM-inspired spatial attention (FSA) blocks.
+Layer normalization (LN) before the GA and after the summing block is provided to stabilize the training process by reducing internal covariate shifts.
+Dropout after the GA block is included for the weight regularization.
+The complete forward model, along with the detailed diagram of FSA is shown in Fig.
+9.
+Each FSA consists of an input convolutional layer with a different kernel size followed by a spatial feature attention with learnable weights.
+The kernel sizes of the input convolutional layer in different FSA heads are selected as $\alpha=1,3,$ and $5$ depending on the complexity of the input geometry.
+If $x_{f}=f^{\alpha\times\alpha}(x)$ is the output of the input convolution, the geometrical feature output based on neighborhood dependencies is obtained using the following equation (1).
+$x_{a}^{s}=f^{\beta\times\beta}\left(\left[w_{1}F^{s}_{avg}(x_{f}),\;\;w_{2}F^{s}_{max}(x_{f})\right]\right)\implies x_{o}=x_{f}\otimes\sigma\left(x_{a}^{s}\right)$ (1)
+where, $f^{\beta\times\beta}(.)$ denotes the convolution with kernel size $\beta$, $\sigma(.)$ denotes the sigmoid operation, and $\otimes$ represents the matrix multiplication.
+$w_{1}$ and $w_{2}$ are learnable weight parameters, initialized with unity.
+The kernel size $\beta$ is a constant value equal to 7 for our case.
+Learnable weights in the FSA block are optimized to learn the inter-dependencies of adjacent connections in the spatial dimension.
+Overall, the whole process attempts to mimic the finite element model to map the vibration spectrum with the metastructure geometry.
+[…]
+The output of the GFE module is processed by a regression head that transforms the geometric features to the transmissibility of dimension $(1000,1)$.
+The fully connected (FC) layers in the regression head are associated with Parametric Rectified Linear Units (PReLU) activation.
+The PReLU activation function facilitates learnable non-linearity to map the highly non-linear nature of the transmissibility.
+The second FC layer is provided with material properties along with geometric features.
+These material properties are modulus of elasticity $(E)$, Density $(\rho)$, Poisson’s ratio $(\nu)$, and material damping $(d)$.
+The output of the last-second FC layer is applied to layer normalization (LN) to stabilize the learning.
+Finally, the last FC layer with linear activation converts the features to the transmissibility having dimension (1000, 1)10001(1000,\;1)( 1000 , 1 ).
+Figure 10: Detailed block diagram of the inverse design model.
+(a) Complete block diagram of inverse design model transforming input transmissibility to metastructure geometry.
+(b) Detailed block diagram of spectrum self-attention responsible for capturing the sequential dependencies of the transmissibility and generating a rectified latent space.
+The spectrum position encoding is specially designed using the Gaussian function to emphasize the specific frequency range in the non-linear spectrum.
+(c) Detailed block diagram of the multiscale residual network.
+The kernel sizes of the convolutional blocks in three channels are $(1\times 1)$, $(3\times 3)$, and $(5\times 5)$, selected to transform the latent space into the three levels of interlaced structural geometry.
+4.2 Inverse Design Model
+The inverse design model is developed to obtain the metastructure geometry using transmissibility data for the desired bandgap.
+The key constituents of the inverse design model are Gaussian self-attention to convert transmissibility to latent space, down sampling network, multiscale residual network to produce interlaced complex structure, and finally upsampling through transpose 2D convolution as shown in Fig 10(a).
+The Gaussian self-attention is a multi-head self-attention mechanism with a Gaussian function for 1D spectrum position encoding.
+It captures the sequential dependencies of the transmissibility and generates a rectified latent space of shape $(3\times 128\times 256)$, shown in Fig 10(b).
+Since the transmissibility is the amplitude versus frequency spectrum, a position encoding must be added to represent the absolute position of the amplitude against the log-scale frequency bin.
+Therefore, a learnable position encoding is defined using sine and cosine of Gaussian functions of log-scale frequency bin to represent the even and odd positions, shown in Equations (2) and (3), respectively.
+$\text{PE}(2p)$ $=$ $\sin\left(\exp\left({-\frac{(\log(p)-\mu_{0})^{2}}{2\sigma_{0}^{2}}}\right)\right)$ (2)
+$\text{PE}(2p+1)$ $=$ $\cos\left(\exp\left({-\frac{(\log(p)-\mu_{0})^{2}}{2\sigma_{0}^{2}}}\right)\right)$ (3)
+where $p$ = position index, $\mu_{0}$ and $\sigma_{0}$ are the learnable parameters for the dynamic allocation of the mean and standard deviations of the Gaussian distribution for 1D position encoding.
+$\mu_{0}$ is initialized with half of the sequence length, whereas $\sigma_{0}$ is initialized with unity to ensure small spread.
+The optimization of these parameters enables the Gaussian function to emphasize the specific frequency range in the non-linear spectrum.
+The latent space generated by the spectrum self-attention is further extended to a new higher space using reflection padding with a reflection of 3 pixels and then non-linearly down-sampled back to the same dimension using multiple consecutive convolutional layers with ReLU activation.
+Now the transformed latent space is passed through a multiscale residual network having three parallel convolutional paths with three different kernel sizes: $(1\times 1)$, $(3\times 3)$, and $(5\times 5)$, shown in Fig.
+10(c).
 These sizes have been selected based on the institution to transform the latent space into the three levels of interlaced structural geometry.
@@
 The final output of the multi-scale residual network is the interlaced patterns.
-It is now upsampled using three layers of 2-D transposed convolution and then transformed non-linearly using 2-D convolution with hyperbolic tangent (Tanh) activation function to produce the final geometry.
-Training
-For the training of the proposed AI-based inverse design of the metastructures, a dataset having 720 samples of different metastructures is prepared using different interlaced structures as shown in supplementary Fig.
-S12.
-In the unit cell of these structures, the primary structure is a honeycomb and the secondary structures are different combinations of honeycomb, star, circle, hierarchical honeycomb, and auxetic interlaced structure.
-Then the corresponding transmissibility is calculated using FEM analysis.
-The samples are divided into two portions: training and validation with a ratio of 90:10. The AI-based inverse design mechanism is trained in two phases.
+It is now up-sampled using three layers of 2D transposed convolution and then transformed non-linearly using 2D convolution with hyperbolic tangent (Tanh) activation to produce the final geometry.
+4.3 Dataset preparation
+Figure 11: Metastructure of primary geometry as honeycomb and secondary geometry (interlaced structures) including honeycomb, auxetic, star, circle, and hierarchical honeycomb used to design different combinations of metastructures for data set generation using numerical simulation.
+A transmissibility dataset was generated for various unique metastructures, having multiscale geometrical designs.
+These metastructures featured a primary honeycomb structure with different secondary interlaced structures: honeycomb, star, circle, hierarchical honeycomb, and auxetic as shown in Fig.11.
+By integrating these secondary structures with the primary honeycomb, a variety of 720 distinct metastructures having different transmissibilities were created.
+The transmissibility of each structure was calculated by post-processing the frequency response obtained from Finite Element (FE) simulations conducted in Ansys.
+Therefore, the training dataset contains 720 samples of metastructure and corresponding transmissibility.
+4.4 Training
+The proposed frameworks of forward analysis and the inverse design models were trained on the metastructure-vs-transmissibility dataset as described above.
+The total samples in the dataset were divided into two portions: training and validation with a ratio of 90:10, resulting in 648 samples in the training set and 72 samples in the validation set.
+The training of the proposed framework was accomplished in the following two phases.
   1. i)
-    Forward Analysis Model: In the first phase, the forward analysis model is trained using the training samples to predict the transmissibility of a given metastructures.
-The training loop is depicted in Fig.
-7(a).
+    Forward Analysis Model: In the first phase, the forward analysis model was trained using the training samples to predict the transmissibility of a given metastructures.
+The training loop has been depicted in Fig.
+8(a).
+The weight optimization was achieved by minimization of the mean square error between the predicted transmissibility and the true transmissibility.
+The model was trained for 300 epochs using the Nesterov-accelerated Adaptive Moment Estimation (NAdam) as the optimization algorithm with dynamic allocation of learning rate by the Cosine Annealing learning rate scheduler (details can be found in [63] and [64]).
 The comparison of the predicted transmissibility obtained by the forward analysis model and the true transmissibility obtained by the FEM analysis are compared in Fig.
-8.
+12.
   2. ii)
-    Inverse Design Model: Once the forward analysis model is fully trained, it is used as a surrogate model for the training of the inverse design model as displayed in Fig.
-7(b).
-The loss function $(\mathcal{L})$ for the weight update of the inverse design model is constructed as provided in equation (1)
+    Inverse Design Model: The trained forward model, obtained from the first phase of the training was used as a surrogate model for the training of the inverse design model as displayed in Fig.
+8(b).
+The weight optimization was accomplished again using the NAdam optimization algorithm and dynamic allocation of learning rate using the Cosine Annealing learning rate scheduler (details can be found in [63] and [64]).
+A composite loss function $(\mathcal{L})$ for the weight update of the inverse design model is defined in Eq.
+(4)
     $\mathcal{L}=\mathcal{L}_{1}\;+\;\beta_{1}*\mathcal{L}_{2}\;+\;\beta_{2}*\mathcal{L}_{3}$ (4)
-    where, $\beta_{1}$ and $\beta_{2}$ are the penalty terms for the inclusion of losses $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$.
+    where, $\mathcal{L}_{1}$ = mean square error (MSE), $\mathcal{L}_{2}$ = structural similarity loss (SSL), and $\mathcal{L}_{3}$ = maximum mean discrepancy (MMD).
+$\beta_{1}$ and $\beta_{2}$ are the penalty terms for the inclusion of losses $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$.
 The values of $\beta_{1}$ and $\beta_{2}$ were selected as 0.1 and 0.6, respectively.
-    $\mathcal{L}_{1}\;=\;\text{MSE}$: mean square error between predicted transmissibility on generated metastructure and the transmissibility obtained from the FEM method on original metastructure.
-    $\mathcal{L}_{2}\;=\;1-\text{SSIM}$: structural similarity index measure between the generated metastructure and the original metastructure [64].
-    $\mathcal{L}_{3}\;=\;\text{MMD}$: maximum mean discrepancy between the generated metastructure and the original metastructure (MMD) [65].
-    Further details of these losses are provided in the supplementary material, Section E.3.
-Validation of AI-model for metastructure design
-The metastructure produced by the inverse design model contains colored noises due to training with a small value of batch size (small batch size was chosen due to the limited number of samples in the dataset).
-Therefore, a noise removal filter is applied to produce a clear image of the obtained metastructure during validation.
-The algorithm for the noise removal is provided in supplementary material, Section F.
-The AI-based inverse design of metastructures is validated under two cases:
+The loss terms MSE, SSL, and MMD are defined as follows:
+    1.
+Mean square error (MSE): MSE represents the mean deviation of the predicted transmissibility of the generated metastructure and the true transmissibility of the original metastructure.
+If $a_{i}\in\mathcal{A}$ and $\hat{a}_{i}\in\hat{\mathbf{A}}$ are $i^{th}$ data point in the true and predicted transmissibilities, the MSE is defined in Eq.
+(5).
+      $\mathcal{L}_{1}(\mathcal{A},\hat{\mathcal{A}})=\frac{1}{n}\sum_{i=1}^{n}\left(a_{i}-\hat{a_{i}}\right)^{2}$ (5)
+    2.
+Structural similarity loss (SSL): SSL represents the structural misalignment of the generated metastructure and the original metastructure.
+It is defined in terms of structural similarity measure as shown in Eq.
+(6)
+      $\mathcal{L}_{2}(\textbf{X},\hat{\textbf{X}})=1-\frac{\sigma_{\textbf{X}\hat{\textbf{X}}}}{\sigma_{\textbf{X}}\sigma_{\hat{\textbf{X}}}+c_{0}}$ (6)
+      where $\sigma_{\textbf{X}\hat{\textbf{X}}}$ = covariance between input image X and the generated image $\hat{\textbf{X}}$, $\sigma_{\textbf{X}}$ = local variance (standard deviation) of the input image X, $\sigma_{\hat{\textbf{X}}}$ = local variance (standard deviation) of the generated image $\hat{\textbf{X}}$, and $c_{0}$ = small constant for division stabilization.
+The inclusion of SSL with a penalty factor of $\beta_{1}$ helps the algorithm to force the weight update in the direction to attain the structural similarity between the generated metastructure and the original metastructure.
+Since the SSL is the opposite of the structural similarity, therefore, $(1-SSL)$ is used to measure the similarity score of the generated image to the original image [65].
+    3.
+Maximum mean discrepancy (MMD): MMD provides the statistical difference (non-parametric distance) between two probability distributions on reproducing kernel Hilbert space (RKHS) [66, 67].
+Therefore, MMD can be used for comparing the statistical distributions of the two images by mapping them into a higher-dimensional feature space using a kernel function.
+If $x_{i}\in\textbf{X}$ and $\hat{x}_{j}\in\hat{\textbf{X}}$ are $i^{th}$ and $j^{th}$ data points in the original and the generated image space, respectively, then the MMD of the two distribution in RKHS is defined in Eq.
+(7)
+      $\mathcal{L}_{3}(\textbf{X},\hat{\textbf{X}})=\left\|{\frac{1}{{n}}\sum\limits_{i=1}^{{n}}{f\left({{x_{i}}}\right)}-\frac{1}{{\hat{n}}}\sum\limits_{j=1}^{{\hat{n}}}{f\left({{\hat{x}_{j}}}\right)}}\right\|_{\mathcal{H}}$ (7)
+      where, $f(.)$ denotes the kernel function $f:X,X\to{\mathcal{H}}$, ${\mathcal{H}}$ be the universal RKHS, $n$ and $\hat{n}$ be the number of pixel points in the original and the generated image X and $\hat{\textbf{X}}$ respectively.
+The MMD, calculated between the original metastructure and predicted metastructure, with a penalty factor of $\beta_{2}$, forces the training algorithm to minimize the discrepancies of the generated image compared to the original image.
+    The inclusion of SSL and MMD in the loss function helps the training algorithm to optimize the inverse design model to learn the original geometry and makes it less dependent on the performance of the forward analysis model in the tandem network.
 Figure 12: Validation results for AI-based inverse design of metastructure.
@@
 (c) Metastructure predicted (generated) by the inverse design model using the input transmissibility.
-SSIM = 0.991, closer to unity validates the exact match with the original image.
+The similarity score = 0.991, closer to unity validates the exact match with the original image.
 (d) Comparing the transmissibilities of the original metastructure obtained from the forward model and FEM.
-(e) Comparing the transmissibilities of the generated metastructure obtained from the forward model and FEM.
-Validation on known metastructure as test sample:
-The comparison of generated metastructure for one test sample from the test data (10% of available 720 samples) is provided in Fig.
-8.
+(e) Comparing the transmissibilities of the generated metastructure obtained from the inverse model and FEM.
+4.5 Validation of AI-model for metastructure design
+The metastructure produced by the inverse design model contains colored noises due to training with a small value of batch size (small batch size was chosen due to the limited number of samples in the dataset).
+Therefore, a noise removal filter is applied to produce a clear image of the obtained metastructure during validation.
+The algorithm for the noise removal is provided in supplementary material, Section E.
+The AI-based inverse design of metastructures is validated under two cases:
+4.5.1 Validation on test samples of metastructure:
+The comparison of generated metastructure for one test sample from the test dataset (10% of available 720 samples) is provided in Fig.
+12.
 Figures 12(a), (b) show the test metastructure and its transmissibility obtained by the FEM method which is assumed to be true (input) transmissibility for inverse design.
-The predicted metastructure by the inverse design model exactly matches the original metastructure with an SSIM score of 0.991 as shown in Figure 8(c).
+The predicted metastructure by the inverse design model exactly matches the original metastructure with the similarity score of 0.991 as shown in Figure 12(c).
 Forward analysis is performed on both the original and the predicted metastructures.
 The predicted transmissibilities are compared with the input (true) transmissibility in Fig.
-8(d) and (e).
+12(d) and (e).
 It can be observed that MSEs for transmissibility obtained from the original metastructure and the predicted metastructure are 0.0834 and 0.1939 respectively.
-The MSE for predicted transmissibility by forward analysis on the predicted metastructure is comparatively larger because of the existence of noises in the predicted metastructure.
-Validation on Artificial Transmissibility:
+The MSE for predicted transmissibility by the forward analysis on the predicted metastructure is comparatively larger because of the existence of noises in the predicted metastructure.
+Figure 13: Additional validation results on multiscale metastructures with different combinations of the unit cells (Honeycomb as a primary unit cell and auxetic, star, hierarchical honeycomb interlacing as secondary and tertiary).
+To further demonstrate the performance of the inverse design framework, we have provided comparisons of the generated metastructure with the original metastructure having different combinations of unit cells in Fig.
+13.
+The first column of Fig.
+13(a) has honeycomb and star interlacing, Fig.
+13(b) has honeycomb, star, and auxetic interlacing, Fig.
+13(c) has auxetic, circle, honeycomb, and hierarchical honeycomb interlacing with the similarity score of 0.992, 0.992, and 0.990 respectively.
+The third column of Fig.
+13 shows the comparisons of predicted transmissibility of the generated metastructure by the forward analysis model and the true transmissibility evaluated by the FEM on the original metastructure with the MSE of 0.0669, 0.0763, and 0.0741 in respective cases.
+4.5.2 Validation on tailored transmissibility:
 The AI-based inverse design model is further validated for practical deployment to design a metastructure facilitating the formation of a desired frequency bandgap.
-Hence, an artificial transmissibility is created having the desired bandgap and fed to the inverse design model.
-The generated metastructure and the corresponding transmissibilities obtained by the forward analysis model and FEM are shown in Fig.
-9.
-The obtained metastructure is the combination of a honeycomb and hierarchical honeycomb substructures.
-The MSE between the predicted transmissibility and the input transmissibility is 0.7408.
-Figure 9: AI-based inverse design of metastructure for a desired bandgap.
-(a) Artificial (input) transmissibility having the desired bandgap as highlighted.
-(b) The generated metastructure obtained from the inverse design model, which facilitates the desired bandgap.
-(c) Validation of the generated metastructure design by comparing the transmissibilities obtained by the forward analysis model and FEM.
-Discussion
+Hence, an artificial transmissibility is created with the desired bandgap and fed to the inverse design model.
+The generated metastructure and the corresponding transmissibilities obtained by the forward analysis model and FEM analysis are shown in Fig.
+14.
+The first column in Fig.
+14 shows the artificial spectrum created for the tailored bandgap, and the second column presents the metastructures generated from the inverse design model for the input transmissibility with tailored bandgap.
+The third column compares the predicted transmissibility by the forward model and the true transmissibility evaluated by the FEM on the generated metastructure.
+Figure 14: AI-based inverse design of metastructure for tailored bandgaps.
+(a) Transmissibility having tailored bandgap at high frequency.
+(b) Transmissibility having tailored wider bandgap.
+(c) Transmissibility having multiple tailored bandgaps ranging from low frequency to high frequency.
+The predicted metastructure design is the combination of a honeycomb and hierarchical honeycomb substructures as shown in Fig.
+14(a).
+As the attenuation band is in the higher frequency region, hence the inverse design model generates a hierarchical substructure in the generated metamaterial design as previously shown in Fig.
+5(a).
+The MSE between the predicted transmissibility and the input transmissibility is 0.0958.
+Figures 14 (b) and (c) have interlaced structures such as honeycomb, auxetic, circle, and star in predicted metastructure design.
+The tailored transmissibility has a wider bandgap and multiple bandgaps in figures 14 (b) and (c) respectively.
+Hence the inverse design model generated auxetic substructures in the generated metamaterial design.
+5 Discussion
 It has been found experimentally and analytically from the previous study that a simple lattice structure has the ability to block waves for a certain frequency range when it passes longitudinally throughout the lattice.
@@
 Some significant negative peaks are obtained from the transmissibility plots that signify wave attenuation and show good agreement with the simulation results, as shown in Fig.
-5.
-The discrepancy in some amplitude and peak are due to nonconsideration of damping in simulation.
-Apart from damping, printing speed, orientation, infill pattern, and infill density of 3D printing are also affected as discussed and shown in Supplementary Table S2.
+7.
+The discrepancy in some amplitudes and peaks is due to the nonconsideration of structural damping in the simulation.
+Apart from damping, printing speed, orientation, infill pattern, and infill density of 3D printing are also affected.
 These findings unveil a deep connection between lattices and band gaps and have broad potential applications in the field of vibration isolation for high-speed trains and in serving as surface waveguides and absorbers.
 The second part of this study on the inverse design of metastructure using geometric attention-based deep architecture is a benchmark in the field of inverse design of complex structure whether it is metastructure for the desired mechanical response or any other complex pattern for desired physical properties.
-The major challenge in the training of an AI-based inverse design model is the availability of sufficient number of training samples.
+The major challenge in the training of an AI-based inverse design model is the availability of a sufficient number of training samples.
 The higher the complexity of the structure, the higher the number of samples required for the training.
@@
 In summary, our paper reflects the following novelties:
-  • Discovery of novel functionality in a set of ancient patterns from the mechanical metastructure perspective such as vibration transmission and attenuation.
+  • 1.
     Identification of two unique parameters: interlaced metastructure and fractal order in such patterns followed by their application in 3D printed 2D mechanical metastructure panels.
+  • 2.
     Further modifications of such metastructures by including local resonators in the form of mass inserts to improve bandgap characteristics through experimental studies.
-  • Creation of novel AI-based forward analysis model using multiscale geometric attention mechanism which can predict the transmissibility as well as bandgap of a metastructure subjected to base excitation.
-[…]
-  • Further, we have been able to create complex multiscale patterns successfully and predict their bandgap characteristics through the transmissibility plot using the current AI model.
+  • 3.
+    Creation of novel forward analysis model with a multi-head FEM-inspired spatial attention (FSA) to predict the transmissibility of a metastructure subjected to base excitation.
+The Multihead FSA efficiently captures the geometrical dependency of vibration mitigation at different scales of structural complexity.
+  • 4.
+    The newly developed model is highly efficient as it is trained only on 648 training samples.
+This results in predicting the transmissibility with the mean square error of 0.08 while in similar literature, ML models were generally trained on 5000 to 40000 data sets to achieve a similar level of accuracy.
+  • 5.
+    Development of a novel MGSA-based inverse design model with 1D spectrum position encoding and a multiscale residual network to produce complex metastructures, given the transmissibility spectrum for a desired frequency bandgap.
+The proposed inverse design model is found to be much more efficient for the reconstruction of complex metastructure with a similarity score of 0.991.
+  • 6.
+    Further, we have been able to successfully demonstrate the effectiveness of the proposed inverse design model to create complex multiscale metastructures by providing artificial transmissibility for a desired frequency bandgap.
 We believe this will facilitate the generation of many customizable complex patterns for mechanical metastructures efficiently.
-  • A novel inverse design model of metastructure based on NLP inspired multiheaded attention mechanism with 1D position encoding and multiscale residual network has also been demonstrated by providing artificial transmissibility for a desired frequency bandgap.
-The inverse design model is found to be much more efficient to reconstruct the complex metastructure with similarity index of 0.991.
-Conclusions
-This research study highlights the translational application potential of vast range of patterns and motifs developed in the field of architecture into the metamaterial-based wave attenuation.
-Thus, the utilitarian values of aesthetic constructs in the field of vibration and sound engineering may open up new systems for wave control including bandwidth control, subwavelength acoustic imaging, and generation of acoustic black holes.
-Additionally, a systematic investigation process has been created on the basis of root lattice identification, multi-scale lattice formation through a hierarchic system, development of hybrid lattice structure, and the effect of interlacing in the vibration mitigation potential initiated from the historic geometric patterns found in the tomb of I’timad-ud-Daulla.
+6 Conclusions
+This research study highlights the translational application potential of the vast range of patterns and motifs developed in the field of architecture into metamaterial-based wave attenuation.
+Thus, the utilitarian values of aesthetic constructs in the field of vibration and sound engineering may open up new systems for wave control, including bandwidth control, sub-wavelength acoustic imaging, and generation of acoustic black holes.
+Additionally, a systematic investigation process has been created on the basis of root lattice identification, multi-scale lattice formation through a hierarchic system, development of hybrid lattice structure, and the effect of interlacing in the vibration mitigation potential.
 In general, the simple lattice structures typically act as attenuators, however, this study demonstrates that the interlaced structure offers an enhanced vibration attenuation property depending on the interlace geometry with improved control over the attenuation frequency range.
-Such effects can be further enhanced, in the presence of and through strategic positioning of the metallic inserts inside the lattice structure which may result in significant expansion of the bandwidth.
+Such effects can be further enhanced in the presence of and through the strategic positioning of the metallic inserts inside the lattice structure, which may result in a significant expansion of the bandwidth.
 This method has been enhanced with the help of AI-based inverse design models to predict metastructures that exhibit the desired bandgap and the corresponding transmissibility.
@@
 Another possibility of future expansion is related to the insertion of smart materials in the lattice structure such that the stiffness and the damping can be controlled over time; thus, transforming it into 4-D metamaterial.
-Methods
-Experimental testing
-The study utilizes a non-contact technique to measure vibration to assess the response of different latticed metastructures.
-Nine 3D printed samples, each showing unique metastructure, are subjected to dynamic testing using a 3D Laser Doppler vibrometer (LDV) from Polytec (Details of the experimental setup presented in Supplementary Fig.
-S10).
-These samples are attached to an Electrodynamic shaker system in cantilever mode.
-The laser head of LDV captures surface velocity by analyzing the phase shift between the incident and reflected laser beams.
-Data acquisition and signal processing are executed through the NI-DAQ system.
-The base excitation technique is utilized to determine transmissibility.
-In this method, the LDS shaker generates 3200 Fast Fourier Transform (FFT) lines, and the pseudo-random excitation with a frequency range of 0 to 10 KHz.
-Displacements of the top and bottom surfaces of the samples are recorded to calculate transmissibility after post-processing of the data.
-Finite element simulations
-Finite element analysis-based simulation has been carried out using Ansys, employing the Ansys Parametric Design Language (APDL) solver for all the nine structures, and the host material is chosen as PLA (detailed properties given in Supplementary Table S3).
-The entire structure was discretized by using tetrahedron elements with a 2 mm element size, to ensure convergence for interlaced geometries.
-The smaller edges of the base structure were fixed, while the longer edges remained free.
-A dynamic force of amplitude 10 N was applied to the base, and the response was calculated on the opposite phase.
-Modal analysis was performed to identify all modes up to 10 KHz frequency, followed by harmonic analysis to obtain the transmissibility.
-The mode superposition method was used to link the modal analysis results with the harmonic analysis input.
-Frequency response functions (FRFs) were generated over the frequency range of 0 to 10 KHz with a 10 Hz frequency resolution; this step size is preferable to accurately capture the response.
-A structural damping ratio of 0.001 was incorporated to account for the viscoelastic effect of PLA substrate.
 Data and code availability
@@
 The photo credits for Figure 2(a-f) are given to Chitralekha Bhattacharya and Bipasana Bhattacharya.
-Author information
-Authors and Affiliations
-Department of Mechanical Engineering, Indian Institute of Technology Kanpur, India
-Bishakh Bhattacharya, Tanuj Gupta, Arun Kumar Sharma, Ankur Dwivedi, Subhadeep Sahana, Ashish Awasthi
-Department of Precision and Microsystems Engineering, Delft University of Technology, Netherland
-Vivek Gupta
-Department of Mechanical Engineering, Motilal Nehru National Institute of Technology Allahabad, India
-Suryansh Pathak
 Author contribution
-Bishakh Bhattacharya: Conceived the primary idea, reviewed the manuscript, critically examined the results, checked the data, wrote the manuscript, and acquired funding.
-Tanuj Gupta: Expanded the idea jointly, modeled all the structures, and designed and performed numerical simulations.
-Fabricated the samples and carried out related experiments and material testing, documented the results, reviewed the manuscript, critically examined the results, and checked the data, wrote the manuscript.
-Arun Kumar Sharma: Concived the idea of AI-based inverse design of metastructures, and implemented a new strategy of feature learning using multiscale Geometric Attention and the idea of reproducing the geometry with multiple levels of complexity using the multiscale residual network.
+Tanuj Gupta: Conceived and expanded the idea jointly, modeled all the structures, and designed and performed numerical simulations.
+Fabricated the samples by 3D printing and carried out related experiments and material testing, documented the results, reviewed the manuscript, critically examined the results, and checked the data, wrote the manuscript.
+Arun Kumar Sharma: Concived the idea of AI-based inverse design of metastructures, and implemented a new strategy of feature learning using multiscale GA and the idea of reproducing the geometry with multiple levels of complexity using the multiscale residual network.
 Code Implementation, training, and validation of the AI models and writing the AI-based inverse design in the manuscript.
 Critically examined the manuscript.
-Ankur Dwivedi: Conceived and expanded the idea jointly in the inverse design of metastructures, critically examined the results, checked the data, wrote and reviewed the manuscript.
+Ankur Dwivedi: Conceived and expanded the idea jointly for the inverse design of metastructures, critically examined the results, checked the data, wrote and reviewed the manuscript.
 Vivek Gupta: Expanded the idea jointly, reviewed the manuscript, critically examined the results, checked the data, and wrote the manuscript.
@@
 Ashish Awasthi: Carried out numerical simulations to generate synthetic data for training the AI model.
+Bishakh Bhattacharya: Conceived the primary idea, reviewed the manuscript, critically examined the results, checked the data, wrote the manuscript, and acquired funding.
 Competing interest
 The authors declare no competing interests.
-Supplementary Material
-Appendix 1 Supplementary Methods
-A Sample Fabrication
-Figure S1: This figure represents a typical interlaced self-similar lattice metastructure employed for sample fabrication.
-The important geometric parameters are prominently displayed, alongside the designated active lattice region.
-This intricate design serves as the blueprint for all samples of our study.
-Some of the sublayer details are shown in Figures S2 and S3.
-A.1 Structural design
-The modeling of all lattice structures is carried out using SolidWorks modeling software.
-For all the structures, the length ($l$) is taken as 174 mm, width ($w$) 58 mm, and thickness ($t$) 5 mm respectively.
-On both sides of the structure, an extension of a 15 mm solid structure is also included for holding and testing purposes as shown in Fig.
-S1.
-The final dimension of the latticed structure ($L\times b\times t$) is $204\times 58\times 5$ mm which is uniform for all the structures so that the results will not get affected by the size of the samples [66].
-Each sample unit cell has an interlaced structure.
-Some of them are interlaced inside a honeycomb structure, while others are interlaced with different structures like ring or fishtail patterns.
-The unit cell of each structure is shown in Fig.
-S2 and Fig.
-S3 and their detailed dimensions are provided in Table S1.
-Figure S2: Intricate detail of unit cell for each lattice structure (LS1 - LS5). Precise dimensions are provided in table S1 which govern the macroscopic properties of the materials.
-Figure S3: Intricate detail of unit cell for each lattice structure (LS6 - LS9). Precise dimensions are provided in table S1 which govern the macroscopic properties of the materials.
-Table S1: Detailed dimensions of structures (LS1-LS9) represented in Figures S2, S3.
-Parameter Value Parameter Value Parameter Value
-LS1 $l_{1}$ 21 mm LS3 $d_{2}$ 2.33 mm LS6 $r_{2}$ 4 mm
-$t_{1}$ 2 mm LS4 $l_{1}$ 12.89 mm LS7 $l$ 5.1962 mm
-$l_{2}$ 8 mm $l_{2}$ 3.71 mm $t$ 1.5 mm
-$t_{2}$ 1 mm $l_{3}$ 2.71 mm $b$ 0.8 mm
-LS2 $l$ 21 mm $t_{1}$ 2 mm LS8 $l_{1}$ 18 mm
-$t_{1}$ 2 mm $t_{2}$ 1 mm $l_{2}$ 16.17 mm
-$t_{2}$ 1 mm $b$ 6.075 mm $t$ 2 mm
-LS3 $l_{1}$ 21 mm $r$ 5.077 mm LS9 $l$ 21 mm
-$t_{1}$ 2 mm LS5 $r$ 8.5 mm $t$ 2 mm
-$\alpha$ $16.28^{\circ}$ $t$ 2 mm $\alpha$ $80^{\circ}$
-$d_{1}$ 5.5 mm LS6 $r_{1}$ 6 mm
-A.2 Additive manufacturing
-Figure S4: (a) Additive manufacturing: Creating intricate three-dimensional samples layer by layer. (b) Finished 3D printed samples.
-Metamaterial structures were modeled using SolidWorks 2022 and subsequently fabricated using a 3D printer that worked on a fused deposition method and has a printing accuracy of $\pm$ 0.1 mm.
-This technology prints samples layer by layer.
-The opted 3D printing specifications included a layer thickness of 0.2 mm and a $50\%$ infill density of the pattern.
-Due to the limitations of 3D printing technology, it was observed that the mechanical properties of the printed material were affected by factors such as layer orientation, infill density, and infill pattern [67, 68].
-To maintain consistency, all sample structures were printed using the same specifications as shown in Table S2, and the corresponding samples are shown in Fig.
-S4.
-The 3D printing material PLA from eSun was utilized to print the sample.
-The filament diameter of the PLA is 1.75 mm.
-The mechanical properties of PLA sample, detailed in Table S3 include density = 1240 $Kg/m^{3}$, Poisson’s ratio = 0.33, and Modulus of Elasticity = 3.5 $GPa$.
-Table S2: A detailed specification of 3D printing process while printing samples.
-Sr. No. Specification in 3D printing Values
-1 3D printing material PLA of 1.75 mm diameter
-2 Layer thickness 0.2 mm
-3 Infill pattern Zig-zag
-4 Infill density $50\%$
-5 Printing speed $50$ mm/sec
-6 Print temperature $210^{\circ}C$
-7 Temperature of build plate $60^{\circ}C$
-8 Top and bottom layer thickness 0.8 mm
-A.3 Metallic insert fabrication
-Figure S5: (a) Mild steel hexagon inserts cut from plate by wire EDM process which is used as a resonator within the lattice structure. (b) Assembly of inserts within the lattice structure.
-Metallic inserts, when strategically integrated into the matrix of a mechanical metamaterial, may function as local resonators.
-These components play a pivotal role in manipulating the attenuation bandwidth, enabling the design of structures with extraordinary characteristics unattainable through conventional materials.
-Mild steel hexagonal shape inserts have been chosen for the selected interlaced honeycomb lattice.
-Initially, hexagon modeling is done on SolidWorks 2022 and exported as a DFX file for the wire EDM cutting using Ezeecut NXG wire EDM, which has a tolerance of $\pm$ 0.002 mm and a best surface finish of 1.5-2 $\mu m$.
-These hexagon shapes are then cut from a 5 mm thick plate, as shown in Fig.
-S5, resulting in inserts with an average weight of 1.0693 grams with a standard deviation of 0.0022 grams.
-The inserts function as a local resonator within the lattice structure due to their large density difference with PLA.
-The mechanical properties of the mild steel sample, detailed in Table S3, include density of 7850 $Kg/m^{3}$, Poisson’s ratio of 0.33, and Modulus of Elasticity of 200 GPa.
-B Material Testing
-B.1 X-Ray fluorescence spectrometry (XRF)
-Figure S6: X-ray fluorescence (XRF) spectroscopy setup for identifying the elements in mild steel sample.
-Table S3: Mechanical properties of PLA and Mild Steel, where subscript $p$ for PLA and $m$ for Mild Steel.
-Material Property Value Material Property Value
-$E_{p}$ 3.5 $GPa$ $E_{m}$ 200 $GPa$
-$\nu_{p}$ 0.33 $\nu_{m}$ 0.33
-$\rho_{p}$ 1240 $Kg/m^{3}$ $\rho_{m}$ 7850 $Kg/m^{3}$
-The main objective of XRF is to identify and quantify the elements present in the material (stainless steel hexagonal inserts) without destroying them.
-A $25\times 25\times 5$ mm stainless steel sample of metal plate is prepared and sample is typically cleaned to remove surface contaminants and sample is irradiated with high-energy X-rays.
-The atoms in the sample emit characteristic X-rays in response to the excitation, and the emitted X-rays are detected and analyzed to determine the elements present and their concentrations (as shown in Fig.
-S6).
-The details of the element present in the sample are presented in Table S4.
-Table S4: Component analysis of mild steel by X-ray fluorescence spectroscopy (XRF)
-[…]
-1 Fe 99.4
-2 Mn 0.321
-3 Si 0.128
-4 Al 0.115
-5 Cl 0.0491
-6 S 0.0220
-7 Ca 0.0141
-B.2 Mechanical characterization: Dynamic mechanical behavior of PLA
-Figure S7: Result of mechanical analysis. (a) Variation of damping parameter with temperature for PLA at various frequency ranges. (b) Storage modulus of 3D printed structure.
-The DMA7100 is used to analyze the PLA sample (with zigzag infill pattern) of standard dimensions 40 x 10 x 1 mm.
-The measurement mode is selected as the tensile mode, and the frequency mode is chosen as sinusoidal oscillation, providing a temperature sweep from $25$ to $80^{\circ}C$ with a heating rate of $2^{\circ}C/min$.
-The resulting graph depicts the damping parameter against temperature for four different frequency ranges (1, 2, 5, and 10 Hz), as illustrated in Fig.
-S7a The graph for PLA represents the measure of material damping/energy dissipation concerning temperature, with the peak indicating the glass transition temperature of the material.
-Fig.
-S7b represents the variation of storage modulus with temperature.
-The graph property at low temperatures corresponds to a high-frequency range, as depicted in Fig.
-S7.
-The damping parameter remains uniform up to $50^{\circ}C$, signifying that for high frequencies, the damping of the material is also uniform.
-Therefore, the occurrence of the bandgap in the lattice structure at high-frequency ranges is attributed to the lattice geometry rather than the material itself.
-C Finite Element Analysis (FEA) Simulations
-Figure S8: Comparison of properties for fractal (a) Bandgap range comparison for different order (n=0,1,2) of fractal structure (b) Force-deflection relation for given fractals.
-Figure S9: Comparison of (a) antysymmetric, and (b) symmetric structure and their respective FRF plots.
-FEA simulation was performed using Ansys software, focusing on modal and harmonic response modules.
-The primary preparation of the model involved material selection (PLA) and mesh definition.
-In the structural modeling, tetrahedron mesh size of 2 mm was chosen based on the geometry’s size and complexity.
-The mode superposition method links the modal analysis results with the harmonic analysis input.
-Fixed edge boundary conditions were applied to the face receiving the vibration waves.
-In the modal analysis, 100 modes were extracted to obtain a broader spectrum of input frequencies for the harmonic response analysis, leading to more accurate results.
-A dynamic force of amplitude 10 N and frequency varying from 0 to 10 KHz was applied to the same face for vibrational excitation and a damping factor of 0.001 was incorporated for practical considerations.
-C.1 Effect of fractal order and lattice symmetry on bandgap
-Fractal structures represent self-similar properties, meaning their complex geometry repeats at progressively smaller scales.
-These self-similar features have no inherent limit and become increasingly intricate with higher fractal orders.
-Figure S8a illustrates the comparison between different fractal orders and the resulting effect on the bandgap frequency range.
-Figure S8 shows how the base geometry transitions from a simple honeycomb structure (order 0) to a first order structure with three repeating honeycombs to a second order structure [69, 70].
-As the fractal order increases, the bandgap shifts towards higher frequencies, and multiple bandgaps also appear.
-In addition, the stiffness of the structure also increases with increasing fractal order.
-Figure S9 represents the effect of symmetric and antisymmetric structures on wave transmission.
-To investigate this effect, we have created two structures, one with antisymmetry in both longitudinal and transverse directions, and the other with symmetry in both directions.
-It clearly demonstrates that an antisymmetric structure is more effective for vibration isolation applications.
-Apart from antisymmetry, by increasing the waviness of the structure, the bandgap also increases.
-The non-dimensional parameter, Gap-mid ratio ($BG$ ), serves as an expression for the width of the bandgap and the center frequency.
-It provides a measure of the relative size of the bandgap with respect to the band structure.
-A higher gap-mid-ratio indicates a larger and potentially more useful bandgap.
-$BG\%\ =2\frac{(f_{u}-f_{l})}{(f_{u}+f_{l})}\%$ (2)
-where $f_{l}$ and $f_{u}$ are the lower and upper-frequency limit of the band gap.
-D Experimental Method
-Figure S10: (a) The experimental setup, accompanied by the Laser Doppler Vibrometer (LDV), is depicted in the schematic diagram.
-(b) 3D laser head with the controller is featured.
-(c) Power amplifier (d) Electro-dynamic shaker which is employed in setup (e)-(f) Excitation is provided from the base through an electrodynamic shaker, while response is obtained by Laser Doppler Vibrometer focused on the top surface of the work-piece.
-The study utilizes a non-contact technique to measure vibration to assess the response of different latticed metastructures.
-Nine 3D printed samples, each showing unique metastructure, are subjected to dynamic testing using a 3D Laser Doppler vibrometer (LDV) from Polytec (Details of the experimental setup presented in Fig.
-S10).
-3D printed structures are made of PLA (the detailed properties of the material are given in Table S3).
-The properties and composition of the hexagonal shape inserts made of mild steel are shown in Tables S3 and S4.
-These samples are attached to an Electrodynamic shaker system in cantilever mode.
-The laser head of LDV captures surface velocity by analyzing the phase shift between the incident and reflected laser beams.
-Data acquisition and signal processing are executed through the NI-DAQ system.
-The base excitation technique is utilized to determine transmissibility.
-In this method, the LDS shaker generates 3200 Fast Fourier Transform (FFT) lines, and the pseudo-random excitation with a frequency range of 0 to 10 KHz.
-Displacements of the top and bottom surfaces of the samples are recorded to calculate transmissibility after post-processing of the data.
-A retro-reflective tape is also attached to the surface where data has been recorded to reduce the scattering of the laser beam from the surface and three closely spaced points were taken on both the top and bottom surface for averaging of data to reduce error.
-D.1 Damping characterization of interlaced lattice metastructures using half-power bandwidth method
-Figure S11: A comparison of damping parameters obtained from dynamic testing performed under base excitation in LDV for different lattice structures (LS1-LS9).
-Table S5: The damping value $\zeta$ obtain from dynamic testing performed in LDV.
-Lattice No. -3dB Points (Hz) Damping ratio ($\zeta$) Quality factor (Q)
-LS1 816.7 828.8 0.007386 67.69
-LS2 896.9 903.4 0.003642 137.30
-LS3 1083.4 1093.1 0.004481 111.59
-LS4 1062.3 1075.0 0.005951 84.02
-LS5 785.5 797.0 0.00732 68.30
-LS6 855.2 865.0 0.005654 88.43
-LS7 876.8 896.4 0.005434 92.01
-LS8 738.0 744.5 0.004415 113.25
-LS9 1191.8 1196.2 0.001867 267.82
-Damping in mechanical systems holds significant importance and is represented in numerous formats.
-The most common terms are the amplification factor, also known as the quality factor ($Q$), and the viscous damping ratio often expressed as the fraction of critical damping ($\zeta$).
-The damping parameter was calculated using the half-power bandwidth method based on the damping test results obtained from Laser Doppler Vibrometry (LDV).
-The test utilized 3200 Fast Fourier Transform (FFT) lines and was driven by a shaker with a unity gain setting for each sample.
-Data acquisition and signal processing were performed using NI-DAQ.
-The highest natural frequency is observed for honeycomb-auxetic (LS9) metastructure (1193.75 Hz) and the lowest natural frequency is observed for ring interlaced (LS5) metastructure (787.5 Hz).
-Using the half-power bandwidth method, the damping for each parameter was calculated as shown in Fig.
-S11.
-The highest and lowest damping observed for honeycomb-honeycomb structure (LS1) and honeycomb auxetic (LS9) metastructure are $\zeta=0.007386$ and $\zeta=0.001867$ respectively (as shown in Table S5).
-The formula relates two parameters:
-$Q=\frac{1}{2\zeta}$ (3)
-The Q value is equivalent to the magnitude of the peak transfer function observed in a single degree of freedom subjected to base excitation at its natural frequency.
-In this context, we introduce the quality factor or Q factor, which characterizes the system’s bandwidth in relation to the resonant frequency.
-A low Q factor reduces the significance of the resonant peak, which poses challenges in measuring changes in resonant frequency.
-Alternatively, the Q factor is defined as the ratio of the energy stored in the system to the energy dissipated per cycle.
-A higher Q factor (L9) leads to a lower rate of energy loss, resulting in a sharper peak of resonant frequency.
-$2\zeta=\frac{\omega_{2}-\omega_{1}}{\omega_{n}}$ (4)
-E AI-driven Inverse Design of Metastructures
-Figure S12: Metastructure of primary geometry as honeycomb and secondary geometry has honeycomb, auxetic, star, circle, and hierarchical honeycomb used to design different combinations of metastructures for analysis purpose.
-In Figure S12 we have selected the honeycomb-based metastructure as the primary structure and six different structures as the secondary structure interlaced inside the honeycomb structure.
-With the help of different combinations of secondary structures with primary metastructure, we have generated 720 different metastructures for analysis purposes, out of which one metastructures we can see as a resultant geometry.
-The analysis data has been taken for learning of forward and reverse models.
-E.1 Forward analysis model
-Figure S13: Main block diagram describing the forward analysis model. LN: Layer Normalization, GA: Geometric Attention.
-The Forward analysis block learns the geometrical patterns and their effect on the transmissibility of the metastructure.
-The model is designed to capture geometrical features at multiple levels of hierarchy in terms of the sizes of patterns.
-The block diagram of the forward model is shown in Fig.
-S13.
-The box shown in the red dotted line is the main geometric feature extractor which is repeated N-times to make the model deeper.
-The main constituent of the geometric feature extractor is geometric attention (GA).
-Layer normalization (LN) before the GA and after the summing block is provided to stabilize the training process by reducing internal covariate shifts.
-Therefore, it helps the model to converge faster.
-Dropout after the GA block is included for the weight regularization.
-Geometric Attention (GA): Fig.
-S14(a) shows the schematic diagram of the GA block.
-The GA block contains the FEM-inspired spatial attention (FSA) block to capture the dependencies on adjacent connections in the input geometrical structure.
-To capture the interlaced complexity of the metastructure, multiple parallel blocks of FSA with convolutional modules with different kernel sizes are employed.
-The later part of the FSA is the spatial attention module [61].
-The functioning of the spatial block to transform the features for geometry learning is described by Fig S14(b)
-Figure S14: Block diagram of the geometric attention (GA). (a) Detailed block diagram of GA and (b) the FEM-inspired spatial attention (FSA).
-Regression Head: The last block of the forwards analysis model is the regression head, the detailed block diagram is represented in Fig.
-S15.
-The purpose of this unit is to process the geometric features from the geometric feature extractor and produce the transmissibility as output.
-The regression head first transforms the geometric features to represent with reduced dimension.
-Then, it is flattened and processed by two fully connected (FC) layers with Parametric Rectified Linear Units (PReLU) activation.
-The second FC is provided with material properties along with geometric features.
-These material properties are modulus of elasticity $(E)$, Density $(\rho)$, Poisson’s ratio $(\nu)$, and material damping $(d)$.
-The output of the last-second FC layer is applied to layer normalization (LN) to stabilize the learning.
-Finally, the last FC layer with linear activation converts the features to the transmissibility having dimension (1000, 1)10001(1000,\;1)( 1000 , 1 ).
-Figure S15: Detailed block diagram of the regression head of forward analysis model.
-E.2 Inverse design model
-Figure S16 shows the main block diagram of the inverse design network.
-It mainly consists of spectrum self-attention to convert transmissibility to latent space, down sampling network, multiscale residual network to produce interlaced complex structure, and finally upsampling through transpose convolution.
-Reflection padding of 3 pixels on each side is provided to increase the size of the latent space with reflected boundary.
-Details of the spectrum to latent space conversion using self-attention and the multiscale residual network are shown in Fig.
-S17.
-Figure S16: Main block diagram describing the implementation of the inverse design model.
-Spectrum Self-Attention: The transmissibility spectrum is the frequency response that represents the geometrical properties as well as the material properties of the metastructure.
-To represent the input transmissibility into 2D latent space, the transmissibility spectrum is given to a multi-head self-attention mechanism followed by a multilayer perceptron (MLP) as shown in Fig.
-S17(a).
-Since the transmissibility is the amplitude versus frequency spectrum, a position encoding must be added to represent the absolute position of the amplitude against the frequency in the input sequence.
-A position encoding of the dimension equal to the system model dimension is defined using alternate sine and cosine functions to represent the even and the odd positions respectively.
-For more details, readers can refer to “Attention Is All You Need” [71]
-Since the objective is to transform the transmissibility spectrum into high-dimensional latent space, first the spectrum is projected to higher dimensional space (the selected embedding dimension is $(1,2048)$ using a fully connected (FC) linear layer.
-Then, the position term is embedded with position encoding as described above.
-The spectrum with position encoded is now passed through a multi-head self-attention with the head dimension equal to the embedding dimension.
-Now the attention output is non-linearly transformed to a further higher dimension to make it suitable for reshaping to the latent space dimension of $(3,128,256)$.
-The ReLU activation function is for non-linear transformation in the MLP block.
-Figure S17: Detailed block diagram of (a) spectrum self-attention (b) multiscale residual network.
-Multiscale Residual Network: The reconstruction of metastructure from the latent space produced by the multi-head self-attention module needs transformation of the features of the latent space with multiple scale factor to produce the pattern of multiple hierarchies.
-Therefore, a residual network with multi-scale kernel sizes of the convolutional module is implemented.
-The main workflow includes passing the input into different convolutional paths with varying kernel sizes.
-The selected kernel sizes for our implementation are $(1\times 1)$, $(3\times 3)$, and $(5\times 5)$.
-This selection is based on the requirement to capture the metastructure at different hierarchies.
-Each convolutional path consists of two convolutional layers in series with ReLU activation.
-Instance normalization is provided between the two convolutional modules.
-The convolutional path with kernel size $(1\times 1)$ helps the model to capture the fine-grained features to construct the tertiary complexity of the metastructure.
-Similarly, the convolutional path with kernel size $(3\times 3)$ captures the spatial features to construct the secondary geometrical structures.
-The third path with kernel size $(5\times 5)$ helps the model to extract the broader contextual information for reconstructing the main blocks on the metastructure geometry.
-Instance normalization is applied to normalize the feature maps across the spatial locations only, improving the stability and efficiency of the training process.
-Since the outputs of each path have to be finally concatenated to merge the transformed feature, the dimensionality of the final outcome of each path has to be the same.
-Therefore, reflection paddings with suitable padding sizes are used in each path before the convolutional layer to adjust the dimensionality with reflected padding.
-The path with the convolutional kernel of $(1\times 1)$ does not alter the dimensionality of the input space, no padding is provided in this path.
-To have the same output dimensionality in the convolutional path with kernel sizes $(3\times 3)$ and $(5\times 5)$, the two reflection padding have sizes of $1$ and $2$, respectively.
-The residual connections, also known as skip connections (additive connections) are provided after two convolutional operations.
-This preserves the original input signal, aiding gradient flow and addressing vanishing gradient issues.
-Since the residual connections add the original features directly to the transformed features to produce the required gradients in the image pattern, again instance normalization is provided to provide stability in the training process.
-Finally, the outputs from different kernel paths are concatenated to combine multiscale features into a unified representation.
-The concatenated feature representation is now projected back to the same spatial dimensionality by using a convolution-based projection layer to represent the final outcome with the same dimensionality.
-Thus, the resultant output retains both the patterns obtained by multiscale transformation and the residual characteristics of the original input.
-Concatenating outputs from multiple paths may lead to an increase in feature dimensions and the need for an additional projection layer, thus adding the higher complexity of the model.
-however, it helps greatly to capture multiple geometrical complexities in the metastructures.
-E.3 Loss function
-The three loss functions used to define the total loss for the inverse model training are described below.
-Mean Square Error (MSE): The mean square error (MSE) can be calculated by taking the difference between the actual and predicted parameters from the following expression
-$\mathcal{L}_{1}=\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}-\bar{Y_{i}}\right)^{2}$ (5)
-Structural Similarity Index Measure: The structural similarity index measure is a yardstick for comparing two images [64].
-SSIM assesses the similarity between images based on structural information available within the image through the correlation of pixel intensities.
-The SSIM is calculated by considering three components called luminance (l), contrast (c), and structure (s) through the following formula
-$\text{SSIM}=\frac{\left(2\mu_{x}\mu_{y}+C_{1}\right)\left(2\sigma_{xy}+C_{2}\right)}{{\left(\mu_{x}^{2}+\mu_{y}^{2}+C_{1}\right)}{\left(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2}\right)}}$ (6)
-where $x$ and $y$ refer to images being compared.
-$\mu_{x}$ and $\mu_{y}$ are mean intensities of images.
-$\sigma_{xy}$ is covariance of both images.
-$C_{1}$ and $C_{2}$ are constants included to avoid instability.
-The higher the value of SSIM, the better the match between the two images.
-Therefore, the loss to minimize for better learning can be defined as
-$\mathcal{L}_{2}=1\;-\;\text{SSIM}$ (7)
-The loss term $\mathcal{L}_{2}$ is included in the total loss with a penalty factor of $\beta_{1}$, a positive fractional number.
-Maximum Mean Discrepancy (MMD): Maximum mean discrepancy MMD is a statistical method used to calculate the difference between two probability distributions.
-MMD basically gives the measure of the non-parametric distance between the mean of two distributions on reproducing kernel Hilbert space (RKHS) [72].
-Hence, MMD compares the distributions of two sets of data by mapping them into a higher-dimensional feature space using a kernel function, and then finally, it measures the distance between their means in the new feature space.
-If the means are equal, it implies that the distributions are similar; if the means are far away, then the distributions differ from each other.
-For a universal RKHS, MMD asymptotically tends to zero if the probability distribution of the two data patterns is the same.
-If $x_{i}^{s}\in\textbf{X}^{s}$ and $x_{j}^{t}\in\textbf{X}^{t}$ are $i^{th}$ source data and the $j^{th}$ target data in the source space $\chi^{s}$ and the target space $\chi^{t}$ respectively, then the MMD of the two distribution in RKHS is defined from the below expression
-$\mathcal{L}_{3}(\textbf{X}^{s},\textbf{X}^{t})=\left\|{\frac{1}{{n^{s}}}\sum\limits_{i=1}^{{n^{s}}}{f\left({{x_{i}^{s}}}\right)}-\frac{1}{{n^{t}}}\sum\limits_{j=1}^{{n^{t}}}{f\left({{x_{j}^{t}}}\right)}}\right\|_{\mathcal{H}}$ (8)
-where, $f(.)$ denotes the kernel function $f:X,X\to{\mathcal{H}}$, ${\mathcal{H}}$ be the universal RKHS, $n^{s}$ and $n^{t}$ be the number of samples in the source and the target datasets $\textbf{X}^{s}$ and $\textbf{X}^{t}$ respectively.
-For our case, loss due to MMD is calculated between the original metastructure and the predicted metastructure.
-It is included in the total loss with a multiplication factor of $\beta_{2}$, which is a positive fractional number.
-The inclusion of MMD in the total loss term forces the algorithm to minimize the discrepancy
-F Image Enhancement for Noise Removal
-The metastructure generated by the inverse design model contains lots of colored noises.
-These noises must be removed for a clear display of the metastructure geometry.
-The following steps are followed to remove the noises from the generated metastructure.
-  1. i)
-    Convert to Grayscale: Convert the input image $I(x,y)$ to a grayscale image $Gr(x,y)$ using:
-    $Gr(x,y)=0.299R(x,y)+0.587G(x,y)+0.114B(x,y)$ (9)
-    where, $Gr(x,y)$ is the spatial dimension of the grayscale image.
-Similarly, $R(x,y)$, $G(x,y)$, and $B(x,y)$ represent the two-dimensional space of reg, green, and blue channels of the image, respectively.
-  2. ii)
-    Apply mask to isolate the structure: Create a binary mask $M_{\text{gray}}(x,y)$ using a gray color threshold value:
-    $M_{\text{gray}}(x,y)=\begin{cases}255&\text{if }Gr(x,y)<T_{\text{gray}}\\ 0&\text{otherwise}\end{cases}$ (10)
-    Apply the binary mask defined above to extract the gray-colored structural regions of the whole image
-    $S_{\text{gray}}(x,y)=G(x,y)\frac{M_{\text{gray}}(x,y)}{255}$ (11)
-  3. iii)
-    Detect black region: Define the lower and upper thresholds for black regions as follows:
-[…]
-    Also, create a binary mask $M_{\text{black}}(x,y)$ for black regions using:
-    $M_{\text{black}}(x,y)=\begin{cases}255&\text{if }Tb_{\text{lower}}\leq I(x,y)\leq Tb_{\text{upper}}\\ 0&\text{otherwise}\end{cases}$ (13)
-  4. iv)
-    Apply black mask: Apply $M_{\text{black}}(x,y)$ to preserve non-black regions:
-    $R_{\text{non-black}}(x,y)=I(x,y)\frac{M_{\text{black}}(x,y)}{255}$ (14)
-  5. v)
-    Combine the results: Combine $R_{\text{non-black}}(x,y)$ and $S_{\text{gray}}(x,y)$ to produce the final image:
-    $F(x,y)={w}_{1}R_{\text{non-black}}(x,y)+{w}_{2}S_{\text{gray}}(x,y)$ (15)
-    Here, $w_{1}=1$ and $w_{2}=1$ are weight parameters to combine the masked image data.
-  6. vi)
-    Normalization: Finally, normalize $F(x,y)$ to the 8-bit range and display as image:
-    $F(x,y)=255\left(\frac{F(x,y)-F_{\min}}{F_{\max}-F_{\min}}\right)$ (16)
-Appendix 2 Supplementary Discussion
-Additive manufacturing of material provides an edge to print some intricate and complex structures precisely for experimental validation, which is difficult to do with traditional machining processes[73, 74].
-Within the regular lattice, the lattice metamaterial has a complex structural geometry designed to effectively control and manipulate vibration waves that pass through it, influencing the material’s overall properties.
-The interlaced structure gives us an edge by providing more parameters to generate a variety of structures.
-By changing geometric properties, we can make these structures tuneable.
-Additionally, the application of local resonators allows us to tune the bandwidth for more specific regions of interest.
-For effective analysis of the materials, the DMA test of PLA and the XRF test of the mild steel sample has been done.
-For further analysis of the interlaced structure, a self-similar interlaced structure which is also called a fractal, similar FEM analysis has been done.
-These properties are essential to designing vibration isolators for sensitive structures.
-A total of 720 data samples with honeycomb-based interlaced metastructure and corresponding transmissibility spectrum have been prepared to develop an AI-based inverse design of metastructure.
-This development is accomplished in two phases of training of deep learning models.
-In the first phase, the images of metastructure geometry were used to train a forward analysis model to predict the corresponding transmissibility spectrum.
-The forward analysis model works based on geometrical feature extraction using the multi-head GA module.
-The multi-head GA has been designed to learn the hierarchies of the metastructure geometry in terms of multiple kernel sizes of convolutional-based feature transformation.
-The second phase of the training involves the development of an inverse design deep learning model that uses the transmissibility spectrum to produce the corresponding metastructure.
-The inverse design model has been proved to be a breakthrough in designing a metastructure for the desired frequency bandgap.
-Therefore, given the desired frequency bandgap, an artificial transmissibility spectrum can be produced and applied to the inverse design model to generate a suitable metastructure that exhibits the frequency response of the desired bandgap.
-This also opens doors for exciting applications in areas such as vibration damping, energy absorption, and even the creation of materials with programmable mechanical behavior.